### v2

# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

Abstract
Objective: To improve prediction of Chronic Kidney Disease (CKD) progression to End Stage Renal Disease (ESRD) using machine learning (ML) and deep learning (DL) models applied to an integrated clinical and claims dataset of varying observation windows, supported by explainable AI (XAI) to enhance interpretability and reduce bias.
Materials and Methods: We utilized data about 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018. Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows. Feature importance and Shapley value analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, notably for African American patients.
Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities.
Conclusion: This study presents a robust framework for predicting ESRD outcomes in CKD patients, improving clinical decision-making and patient care through multi-sourced, integrated data and AI/ML methods. Future research will expand data integration and explore the application of this framework to other chronic diseases.
1 Introduction
[…]
Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 4 or administrative claims data 5. While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression. For instance, Sharma et al. 6 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al. 5 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database. However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics. For example, Tangri et al. 7 used clinical variables like age, gender, eGFR, and albuminuria to predict ESRD progression, and Keane et al. 8 utilized serum creatinine and urine protein levels for high-risk identification. Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices. Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
In the field of explainable AI (XAI) for healthcare, established methods like SHapley Additive exPlanations (SHAP) 9 and Local Interpretable Model-agnostic Explanations (LIME) 10 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models. However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets. At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus. This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making and ultimately improving patient outcomes.
2 Objective
This research aims to evaluate multiple prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data. Our primary objective is to estimate the probability that a CKD patient will develop ESRD in the future based on the data from a specified observation window prior to the diagnosis of ESRD, represented as follows:
$P\left(\mathrm{ESRD}_{i}=1\mid t,X_{\text{clinical},i}(t),X_{\text{claim},i}(t)\right),$
where $X_{\text{clinical},i}(t)$ and $X_{\text{claim},i}(t)$ denote the clinical and claims data observed up to time $t$ for patient $i$ who progresses to ESRD after the observation window.
We address two key scenarios in our study:
  • Model Robustness: We train and test different analytical models on the same dataset to ensure consistency and comparability, allowing us to select the best-performing model with a focus on robustness and reliability.
  • Clinical Utility: In clinical settings, the likely duration of CKD progression to ESRD is unknown. Our model aims to predict the likelihood of ESRD in the future and suggests an optimal observation window based on the duration of data available on the patient from the time of CKD Stage 3 diagnosis in time steps of 6, 12, 18, 24 and 30 months.
To address these objectives, we focus on identifying the best data-driven observation window while maintaining an acceptable error rate $\epsilon(t)$, represented as follows:
$T_{\text{obs }}=\min\left\{t\mid\epsilon(t)\leq\epsilon^{\prime},t\in\{6,12,18,24,30\}\right\},\vspace{-.5cm}$
where $\epsilon^{\prime}$ is defined as:
$\epsilon^{\prime}=\left\{\begin{array}[]{ll}\textnormal{user-defined error rate}&\textnormal{if specified},\\ \min_{t\in\{6,12,18,24,30\}}\epsilon(t)&\textnormal{otherwise}\end{array}\right.$
Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
3 Materials and Methods
3.1 Data
This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data. The de-identified administrative claims dataset, obtained from a major health insurance provider and approved by their Institutional Review Board, includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. It encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with Chronic Kidney Disease (CKD) and managed by a leading community nephrology practice that shared their de-identified clinical data.
The clinical dataset was sourced from the EHR of the nephrology practice, containing laboratory results, patient demographics, diagnostic details, and medication records from 1998 to 2018. For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data. Usage of the clinical data was approved by the Institutional Review Board of the University.
Preprocessing of the claims data involved removing duplicates, excluding records without any stage of CKD diagnoses, and eliminating entries with negative costs to ensure data integrity. After cleansing, the claims dataset included 5,317,178 individual claims across 7,129 unique patients. The clinical dataset was similarly refined to include 433,421 lab records for 10,326 patients. These datasets were then integrated using unique, de-identified patient IDs, providing a comprehensive view of each patient’s medical and claims history. This integration was crucial for accurately identifying ESRD predictors using multiple analytical models.
3.2 Research Framework
Figure 1: Overview of research framework and training strategy. Part A provides an overview of the data preparation, modeling and validation, and additional analyses pipeline. Part B illustrates the cohort identification process, detailing the inclusion and exclusion criteria applied to claims and clinical records. Part C outlines the modeling strategy.
As illustrated in Fig.1, our framework for predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) consists of three key stages: Data Preparation, Modeling & Validation, and Additional Analyses form robustness and clinical utility.
3.2.1 Data Preparation
Data Pre-processing is the foundation of the framework, involving the integration and thorough cleaning of claims and clinical datasets. This Part A stage includes data preprocessing steps such as cleansing, normalization, missing value imputations and log transformations as appropriate for data analysis.
Cohort Identification: Following this, in Part B, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period. Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
The study cohort was curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists. We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management. Further refinement excluded patients who progressed to ESRD within the observation window, resulting in a final cohort with uninterrupted, detailed medical and claims histories.
Feature Engineering, depicted in Part C, involved identifying and creating predictive variables essential for modeling. With claims data, we categorized the features used for prediction into two main groups. The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision. These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
The second group, comorbidity-based features, includes patient-specific clinical information such as the duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia. These features are crucial for understanding CKD progression and were statistically analyzed for their significance in predicting ESRD outcomes (as detailed in Table 1(a) and Table 1(b)).
Table 1: Summary of Descriptive and Categorical Statistics for CKD Features
(a) Descriptive statistics for claims-driven and clinical-driven features; means and standard deviations (StD) for ESRD vs. non-ESRD populations. The $p$-value is based on the t-test of the difference between the ESRD and non-ESRD populations. A significance level of $0.05$ is used.
Category Feature Name Overall Mean (StD) ESRD Mean (StD) Non-ESRD Mean (StD) $p$-value
Claims-driven Count of pharmacy claims (CM1) 110 (88) 120 (94) 109 (87) 0.035
Count of inpatient claims (CM2) 3.74 (3.80) 3.75 (3.43) 3.74 (3.87) 0.959
Count of outpatient claims (CM3) 23.69 (22.07) 27.78 (28.75) 23.09 (20.83) 0.003
Count of professional claims (CM4) 89.73 (68.53) 105.37 (77.46) 87.43 (68.01) 0.000
Net cost of pharmacy claims (CM5) 10645 (20303) 12053 (17596) 10440 (20662) 0.110
Net cost of inpatient claims (CM6) 30038 (36080) 33909 (53540) 29440 (32541) 0.121
Net cost of outpatient claims (CM7) 8657 (17495) 9354 (17522) 8554 (17492) 0.415
Net cost of professional claims (CM8) 12137 (13707) 15512 (18657) 11640 (12748) 0.000
Range of claims costs (CM9) 9173 (15910) 11352 (32606) 8852 (11550) 0.147
Standard deviation of claims costs (CM10) 766 (879) 831 (1263) 757 (806) 0.277
Clinical-driven Age at CKD stage 3 first diagnosis (CL1) 72.68 (11) 70.13 (10.37) 73.04 (11) 0.000
CKD stage 3 duration time (days) (CL2) 660 (179) 543 (250) 677 (160) 0.000
Number of emergency department visits (CL3) 2.11 (2.94) 2.18 (2.63) 2.01 (2.99) 0.258
(b) Categorical clinical-driven features; proportions of ESRD vs. non-ESRD populations. The difference between the ESRD and non-ESRD populations is statistically significant for most of the clinical-driven categorical features based on chi-squared test of independence results with a significance level of $0.05$.
Category Feature Name ESRD Proportion Non-ESRD Proportion ${\chi}^{2}$ Statistics $p$-value
Clinical-driven Gender (CL4) (male:female) 60:40 52:48 6.73 0.009
Occurrence of CKD stage 4 (CL5) 43:57 12:88 217.8 0.000
Occurrence of CKD stage 5 (CL6) 5:95 1:99 48.8 0.000
Diabetes (CL7) 73:27 59:41 27.3 0.000
Anemia (CL8) 64:36 62:38 0.58 0.445
Metabolic acidosis (CL9) 25:75 18:82 9.41 0.002
Proteinuria (CL10) 13:87 17:83 3.72 0.054
Secondary hyperparathyroidism (CL11) 33:67 18:82 42.9 0.000
Phosphatemia (CL12) 5:95 3:97 3.55 0.06
Atherosclerosis (CL13) 6:94 14:86 18.04 0.000
Heart failure (CL14) 5:95 11:89 13.7 0.000
Stroke (CL15) 1:99 3:97 4.58 0.032
Conduction & dysrhythmias (CL16) 5:95 16:84 29.9 0.000
Hypertension (CL17) 99:1 97:3 3.1 0.078
Clinical data features were categorized into three groups based on consultations with nephrologists, commonly referenced literature, and the characteristics of our datasets. These groups include demographic features, laboratory data features, and comorbidity features.
  • Demographic Features: These include age, gender, ethnicity and BMI, which have been widely used for understanding CKD progression.
  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, bicarbonate, and albumin, which are important for monitoring CKD. Notably, UACR was not selected due to insufficient readings compared to the other features.
  • Additional Comorbidity Features: These capture the presence of additional health conditions such as cardiovascular disease, anemia, and metabolic acidosis, in addition to those identified using claims data, and have been shown to influence CKD outcomes.
3.2.2 Modeling & Validation
To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns. Our modeling strategy (Fig. 1 Part C) involved partitioning data into training (60%), validation (20%), and testing (20%) sets to ensure model reliability. We optimized performance through hyper-parameter tuning and an iterative process to determine the optimal observation window that balances predictive accuracy and timeliness.
Machine Learning Methods
We employed statistical methods, beginning with logistic regression (LR) 11 as a baseline to estimate the probability of ESRD progression using clinical and claims data features. Beyond LR, we utilized ML techniques such as Random Forest (RF) 12 and Extreme Gradient Boosting (XGBoost) 13. RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting. XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
To enhance model interpretability, we conducted cohort-level feature importance analysis to identify key predictors across the dataset. Additionally, SHAP analysis was applied at the individual level to elucidate each feature’s contribution to specific predictions. This dual approach—cohort-level and individual-level analysis—not only provides insights into the drivers of ESRD progression but also improves model transparency, supporting personalized treatment decisions.
Deep Learning Methods
To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 14; 15 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis. This approach captures the temporal evolution of CKD, reflecting its gradual progression and aligning with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data. Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression.
We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 16 for feature extraction, Recurrent Neural Network (RNN) 17, Long Short-Term Memory network (LSTMs 18, Gated Recurrent Unit (GRU) 19 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 20 for sequence modeling. These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision.
Modeling Scenarios: Robustness and Clinical Application
As outlined in the Objective section, our modeling approach addresses two key scenarios:
  • Model Robustness: To ensure model consistency and fair comparability, we train and test on a unified dataset. We address the issue of shrinking cohort size with longer observation windows by selecting cohorts with a minimum of 30 months of data following CKD stage 3 diagnosis. The dataset is partitioned into training (60%), validation (20%), and testing (20%) subsets. By varying training data length from 6 to 30 months within this cohort, we assess model performance under different data availability conditions, thereby evaluating robustness and reliability.
  • Clinical Utility: In clinical practice, the progression timeline from CKD to ESRD is often uncertain. To model this uncertainty, we integrate all time-length cohorts, enabling the prediction of ESRD likelihood and optimal observation windows without precise progression timelines. This mixed-cohort approach enhances model relevance in real-world clinical settings, accommodating variable patient timelines.
3.2.3 Additional Analyses
This stage includes a detailed examination of incorrect predictions, identification of factors contributing to errors, and exploration of potential biases, particularly considering the recent changes to eGFR equations in 2021. These analyses are critical for refining the models and ensuring they deliver fair and accurate predictions.
Analysis of Model Misclassifications
To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set. This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes.
We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error). By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
This detailed error analysis is crucial for improving the model’s accuracy and ensuring its reliability in clinical settings. By understanding the reasons behind incorrect predictions, we can enhance the model’s ability to support clinicians in making more informed and accurate decisions regarding patient care.
Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias. In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare. In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 21, which excluded the race coefficient to improve accuracy and equity 22. This adjustment aimed to enhance prediction precision, especially for minority populations 23; 24; 21.
Our analysis, using data from 2009 to 2018, initially applied the older eGFR equation. To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
4 Results
4.1 Data Integration and Cohort for Analysis
(a) Comparison of CKD stage distribution across claims, clinical, and merged datasets with a 24-month observation window.
(b) Variation in CKD stage distribution as a function of different observation window lengths.
Figure 2: Overview of CKD Stage Distribution in Merged Data Across Observation Window Lengths.
To maximize the utility of multi-sourced integrated data, we merged claims and clinical datasets using unique patient identifiers. The distribution of data for individual and merged datasets is presented in Fig. 2. The analysis cohort consists of patients identified at CKD stage 3, marking the beginning of data collection for each patient. As shown in Fig. 2(a), merging the datasets resulted in a significant reduction in the number of patients—from 5,518 and 6,842 in the claims and clinical datasets, respectively, to 1,422 in the merged dataset with a 24-month observation window. This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis. Fig. 2(b) illustrates the changes in the total number of patients and their distribution across CKD stages with varying observation window lengths. As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis. This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period. Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
4.2 Performance Comparison of Models Across Datasets and Scenarios
4.2.1 Scenario 1: Model Robustness
This scenario focuses on evaluating the robustness and reliability of various models by training and testing them on a consistent dataset. We varied the training data length from 6 to 30 months within a fixed cohort to assess the models’ performance under different data availability conditions. This approach enabled us to evaluate the models’ stability and effectiveness across various observation windows, ensuring consistency and fair comparability.
Table 2: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 1 (Model Robustness). This evaluation focuses on the stability and effectiveness of models across varying observation window lengths. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
(a) Performance of models using claims data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Claims Data Only Modeling Machine Learning Methods Logistic Regression 0.61 0.24 0.64 0.31 0.70 0.32 0.72 0.33 0.71 0.29
Random Forest 0.69 0.26 0.73 0.33 0.74 0.41 0.74 0.36 0.68 0.33
XGBoost 0.70 0.27 0.75 0.35 0.78 0.42 0.75 0.39 0.69 0.38
Deep Learning Methods CNN 0.72 0.36 0.73 0.42 0.81 0.44 0.82 0.45 0.84 0.44
RNN 0.77 0.39 0.75 0.44 0.83 0.49 0.90 0.50 0.82 0.47
LSTM 0.76 0.33 0.79 0.43 0.86 0.45 0.92 0.54 0.87 0.44
GRU 0.70 0.34 0.78 0.46 0.88 0.45 0.91 0.50 0.86 0.42
TCN 0.72 0.36 0.81 0.37 0.85 0.46 0.88 0.52 0.85 0.43
(b) Performance of models using clinical data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Clinical Data Only Modeling Machine Learning Methods Logistic Regression 0.63 0.46 0.70 0.49 0.76 0.51 0.76 0.54 0.75 0.53
Random Forest 0.70 0.53 0.76 0.53 0.80 0.58 0.79 0.58 0.77 0.54
XGBoost 0.73 0.52 0.76 0.55 0.80 0.59 0.80 0.57 0.72 0.55
Deep Learning Methods CNN 0.73 0.51 0.75 0.54 0.80 0.55 0.84 0.56 0.80 0.52
RNN 0.74 0.53 0.77 0.57 0.82 0.60 0.85 0.61 0.79 0.55
LSTM 0.77 0.54 0.78 0.56 0.84 0.63 0.88 0.60 0.82 0.57
GRU 0.80 0.54 0.77 0.57 0.83 0.61 0.87 0.60 0.84 0.53
TCN 0.72 0.50 0.73 0.53 0.80 0.57 0.83 0.61 0.81 0.55
(c) Performance of models using merged data.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Merged Data Modeling Machine Learning Methods Logistic Regression 0.65 0.46 0.69 0.47 0.76 0.54 0.75 0.55 0.78 0.54
Random Forest 0.74 0.44 0.77 0.52 0.83 0.58 0.84 0.60 0.80 0.55
XGBoost 0.73 0.45 0.76 0.50 0.82 0.58 0.85 0.61 0.82 0.59
Deep Learning Methods CNN 0.70 0.45 0.76 0.46 0.77 0.55 0.80 0.56 0.82 0.55
RNN 0.75 0.51 0.80 0.58 0.86 0.60 0.87 0.62 0.85 0.61
LSTM 0.76 0.53 0.84 0.59 0.93 0.63 0.93 0.65 0.84 0.59
GRU 0.75 0.51 0.76 0.56 0.91 0.62 0.90 0.63 0.86 0.57
TCN 0.71 0.49 0.74 0.57 0.88 0.60 0.89 0.61 0.91 0.59
Table 2(c) compares model performance across claims data-only scenario, clinical data-only, and merged data. In claims data-only, ML model XGBoost achieved the highest AUC (0.78) at 18 months, while Random Forest had the best F1 score (0.41). DL based LSTM outperformed all others with an AUC of 0.92 and an F1 score of 0.54 at 24 months. In clinical data-only scenario, Random Forest led with an AUC of 0.80 and an F1 score of 0.58, but LSTM again excelled with an AUC of 0.88 and an F1 score of 0.63. Merged data further improved results, with XGBoost reaching an AUC of 0.85, and LSTM and GRU achieving AUCs of 0.93 and 0.91, with F1 scores of 0.65 and 0.63. Overall, merged data and deep learning methods, particularly LSTM and GRU, delivered the best performance, with the best observation window being 24 months.
4.2.2 Scenario 2: Clinical Application
In this scenario, we addressed the challenge of predicting ESRD development in real-world clinical settings, where the duration of CKD progression is often uncertain. By examining all cohort observation windows, the models predict ESRD likelihood and determine the optimal observation window despite unknown progression timelines. This mixed cohort approach enhances the model’s relevance for clinical practice, accommodating variable patient timelines.
Table 3(c) shows model performance across claims-only, clinical-only, and merged data. In claims-only, XGBoost had the highest AUC (0.76) at 12 months among ML models, while Random Forest had the best F1 score (0.40) at 18 months. However, LSTM outperformed others with an AUC of 0.90 and an F1 score of 0.51 at 24 months. For clinical-only, Random Forest led ML models (AUC 0.79, F1 0.56 at 18 months), but LSTM was superior (AUC 0.85, F1 0.61 at 24 months). Merged data yielded the best results, with XGBoost leading ML methods (AUC 0.84 at 18 months), and LSTM and GRU excelling overall (LSTM: AUC 0.90, F1 0.60; GRU: F1 0.63 at 24 months). Merged data consistently outperformed individual datasets, and deep learning models, especially LSTM and GRU, showed significant performance gains, highlighting the benefits of integrating multiple data sources for accurate clinical predictions.
Comparing the two scenarios, we observe that while both benefit from deep learning models and data integration, they demonstrate different strengths depending on the context. Scenario 1 underscores the importance of stability and consistency across varying data conditions, whereas Scenario 2 highlights adaptability and practical applicability in real-world clinical settings. Together, these scenarios underscore the critical role of deep learning and comprehensive data integration, with 24 months identified as a key observation window that balances robustness and clinical relevance.
Table 3: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 2 (Clinical Application). The model aims to predict the likelihood of ESRD development and suggest an optimal observation window based on data from the patient’s initial stage 3 CKD visit. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
(a) Performance of models using claims data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Claims Data Only Modeling Machine Learning Methods Logistic Regression 0.66 0.22 0.67 0.28 0.70 0.32 0.72 0.34 0.71 0.35
Random Forest 0.72 0.26 0.75 0.30 0.74 0.40 0.73 0.34 0.68 0.31
XGBoost 0.72 0.28 0.76 0.31 0.72 0.36 0.71 0.32 0.69 0.33
Deep Learning Methods CNN 0.70 0.34 0.72 0.48 0.79 0.49 0.88 0.46 0.86 0.42
RNN 0.79 0.43 0.78 0.47 0.82 0.47 0.84 0.49 0.82 0.41
LSTM 0.79 0.33 0.88 0.33 0.84 0.47 0.90 0.51 0.87 0.41
GRU 0.73 0.29 0.85 0.44 0.86 0.50 0.89 0.49 0.86 0.40
TCN 0.72 0.42 0.82 0.38 0.84 0.48 0.87 0.49 0.85 0.40
(b) Performance of models using clinical data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Clinical Data Only Modeling Machine Learning Methods Logistic Regression 0.61 0.45 0.69 0.48 0.78 0.49 0.76 0.50 0.73 0.51
Random Forest 0.72 0.51 0.75 0.53 0.79 0.56 0.73 0.54 0.68 0.50
XGBoost 0.72 0.51 0.76 0.54 0.78 0.58 0.71 0.56 0.69 0.57
Deep Learning Methods CNN 0.70 0.44 0.72 0.46 0.77 0.46 0.80 0.56 0.76 0.54
RNN 0.74 0.48 0.72 0.44 0.78 0.48 0.82 0.59 0.78 0.54
LSTM 0.76 0.44 0.78 0.46 0.80 0.51 0.85 0.61 0.79 0.55
GRU 0.73 0.46 0.78 0.52 0.80 0.50 0.85 0.59 0.77 0.47
TCN 0.69 0.47 0.75 0.48 0.77 0.48 0.81 0.59 0.77 0.51
(c) Performance of models using merged data.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Merged Data Modeling Machine Learning Methods Logistic Regression 0.66 0.44 0.68 0.47 0.74 0.51 0.72 0.51 0.65 0.48
Random Forest 0.72 0.47 0.76 0.51 0.80 0.56 0.81 0.52 0.77 0.53
XGBoost 0.71 0.50 0.79 0.52 0.84 0.57 0.81 0.56 0.79 0.57
Deep Learning Methods CNN 0.60 0.43 0.66 0.43 0.77 0.47 0.78 0.48 0.76 0.50
RNN 0.66 0.49 0.69 0.44 0.80 0.50 0.84 0.55 0.80 0.52
LSTM 0.68 0.49 0.75 0.48 0.89 0.57 0.90 0.60 0.79 0.53
GRU 0.65 0.49 0.76 0.49 0.89 0.55 0.90 0.63 0.79 0.52
TCN 0.65 0.45 0.74 0.50 0.78 0.44 0.88 0.59 0.78 0.53
4.3 Key Factor Analysis Using Explainable AI Techniques
We identified a 24-month observation window as optimal for balancing early detection and prediction accuracy, with LSTM proving most effective among deep learning models and XGBoost leading in machine learning. Focusing on Scenario 2 for its relevance to real-world clinical settings with uncertain CKD progression timelines, we refined our approach further to enhance clinical applicability. We used cohort-level feature importance analysis with XGBoost to determine key predictive features and applied SHAP analysis at the individual patient level to improve model interpretability and transparency.
4.3.1 Cohort-Level Key Feature Identification
Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.  3(a)). The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its significance in predicting ESRD. Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly. The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power. This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice. However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
(a) Top 30 features of importance for the XGBoost model.
(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
Figure 3: Key Feature Analysis: Feature Importance and SHAP Analysis for ESRD Risk Prediction Using A 24-month Observation Window.
4.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach. To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD. Despite both having CKD Stage 5, the contributing features differed significantly, underscoring the need for personalized treatment strategies.
In Figure 3(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue. For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it. For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning. SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile, thereby enhancing the effectiveness of clinical care.
4.4 Analysis of Model Misclassifications: Type I and Type II Errors
To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 4(a). Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered. This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
To understand what drives these misclassifications, we conducted a detailed analysis of Type I and Type II errors, illustrated in Figures 4(c) and 4(b). Type II errors (Figure 4(b)) involve patients wrongly predicted not to develop ESRD but did; we compared these 16 false negatives to the 70 correctly predicted to progress to ESRD. These errors can delay essential treatment, potentially accelerating disease progression. Conversely, Type I errors (Figure 4(c)) involve patients wrongly predicted to develop ESRD who did not. We contrasted 17 false positives with 1,319 correct predictions of non-progression, highlighting how these errors can lead to unnecessary medical interventions, potentially depriving resources from those who truly need them. By examining the feature distributions between these correctly and incorrectly predicted groups, we aimed to identify statistically significant differences that could explain the model’s errors, providing insights for refining the model and improving its predictive accuracy.
(a) Distribution of Predicted Probabilities for Patient Outcome.
(b) Analysis of Type II Errors: Comparison of Patients Predicted Not to Develop ESRD Who Actually Did.
(c) Analysis of Type I Errors: Comparison of Patients Predicted to Develop ESRD Who Actually Did Not.
Figure 4: Analysis of Model Misclassifications: Type I and Type II Errors. The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset. The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
Table 4: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy. This table presents representative subsets of features under various timestamps. Features in bold appear in both analyses.
(a) Type II Error Analysis: Feature Comparison Between False Negatives and True Positives.
Model Mean P-value Significant Timestamp
Correct incorrect
n_claims_I 0.70 0.14 0.048752 Yes 5
n_claims_O 5.71 2.67 0.001325 Yes 2
net_exp_O 3831.83 830.61 0.002747 Yes 0
S5 0.36 0.06 0.001189 Yes 6
S5 0.37 0.49 0.662073 No 7
net_exp_O 4219.01 1044.37 0.001883 Yes 4
(b) Type I Error Analysis: Feature Comparison Between False Positives and True Negatives.
Model Mean P-value Significant Timestamp
Correct incorrect
n_claims_DR 13.52 7.21 0.001156 Yes 3
n_claims_O 3.03 1.69 0.009384 Yes 4
S4 3831.83 830.61 0.002747 Yes 4
S5 0.09 0.35 4.059934e-02 Yes 6
S5 0.08 0.35 3.911497e-02 Yes 7
net_exp_O 1044.82 333.61 0.000324 Yes 4
Based on Table 4(a), CKD Stage 5 emerges as a critical feature. For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, significantly lower than the 0.36 mean for those correctly predicted to progress. This suggests that these patients did not have a Stage 5 record until the subsequent timestamp. The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely. However, after consulting with nephrologists and examining the literature on CKD progression to ESRD, we understand that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
Based on Table 4(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD. This is unusual. One possible explanation could be the issue of data censoring. Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later. Additional data collection beyond 2018 would be needed to update and refine this analysis.
4.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
(a) Incorrect predictive probabilities using the 2009 eGFR equation.
(b) Incorrect predictive probabilities using the 2021 eGFR equation.
(c) Distribution of race, gender, and ESRD status with the 2021 eGFR equation. Categories: Race (0 = White, 1 = Black, 2 = Other), Gender (0 = Male, 1 = Female), ESRD (0 = No ESRD, 1 = ESRD).
(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
Figure 5: Impact of Updated eGFR Equations (2009 vs. 2021) on Prediction Errors and Racial Bias. Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
When comparing the incorrect predictions using the 2009 eGFR equation (Fig  5(a)) with those from the 2021 equation (Fig  5(b)), we observed a decrease in the number of incorrect predictions from 33 to 28. This reduction reflects a slight improvement in the error rate, which decreased from 0.0232 to 0.0228 with the updated 2021 eGFR equation. Despite this improvement, the pattern of incorrect predictions being clustered towards the lower end, as discussed earlier, persists.
To assess whether the updated eGFR equation improves prediction accuracy for minoritized groups, we analyzed the demographic and racial distribution of patients who were incorrectly predicted not to develop ESRD. Fig  5(c) displays the demographic breakdown, including race, gender, and ESRD outcomes, for the newly defined cohort under the 2021 equation. Fig  5(d) reveals changes in incorrect predictions: the number of misclassified white patients increased from 2 to 4, while incorrect predictions for African American patients decreased significantly from 4 to 1. Meanwhile, the count for patients in the ”Other” category rose from 10 to 12.
Given the relatively small sample size for the ”Other” racial category, the increase may not be statistically significant. However, the notable reduction in incorrect predictions for African American patients suggests that the 2021 eGFR equation enhances the model’s prediction accuracy for this group. This improvement is particularly important as it indicates progress towards reducing racial bias in ESRD predictions, aligning with broader efforts to enhance equity in healthcare.
Overall, while the updated 2021 eGFR equation has led to a modest reduction in the overall error rate and appears to improve prediction accuracy for African American patients, further analysis with larger cohorts are needed to assess its impact on other racial groups. This finding underscores the ongoing need to refine prediction models to ensure fair and accurate outcomes across all patient demographics.
5 Discussion
This study underscores the value of integrating patient data from multiple sources, such as claims and EHRs, and of different types, such as clinical comorbidities and utilization costs, combined with deep learning methodologies, particularly in the longitudinal representation of healthcare data, for improving the accuracy of patient outcome predictions. By employing models such as LSTM and GRU, which are adept at capturing temporal dependencies in sequential data, we were able to enhance the predictive power to assess progression to ESRD in CKD patients. The use of a 24-month observation window, identified as the optimal period for accurate predictions, allows for timely and reliable risk assessment, providing a critical balance between early detection and maintaining a manageable error rate. These findings add to the rapidly growing evidence for driving more precise and interpretable predictions for chronic disease management, ultimately supporting better clinical decision-making to improve patient outcomes.
This research bridges a significant gap in risk assessment for chronic disease management studies by integrating clinical and claims data. Previous models have often relied on one data type in isolation, either clinical or claims, which can lead to incomplete representations of a patient’s health journey and the progression of CKD. By combining both data types, our approach offers a more comprehensive view, capturing not only medical indicators but also healthcare utilization patterns, as demonstrated by the features that are identified as most important in ESRD prediction. This integrated data strategy enhances the model’s robustness and applicability across diverse patient populations, addressing socio-economic factors and healthcare interactions that single-source models may overlook. Our findings highlight the importance of a multi-faceted approach to data integration, which can lead to more inclusive and accurate models for managing chronic diseases such as CKD.
A key innovation of our work is the determination of an optimal observation window that minimizes the volume of data and time required for making reliable predictions while maintaining an acceptable error rate. This focus on observation period optimization is crucial for balancing the need for timely interventions with the risks of premature or inaccurate predictions. By rigorously evaluating different observation window lengths, we were able to identify a 24-month period as optimal, providing a critical framework for clinicians to monitor and manage CKD progression effectively. This approach supports a proactive stance in patient care, allowing for earlier detection and management of patients at risk of progressing to ESRD, thereby potentially reducing long-term healthcare costs and improving patient outcomes.
The insights derived from our research have significant practical implications for early intervention strategies in CKD management. By leveraging integrated data and advanced modeling techniques, healthcare providers can better identify patients at high risk of advancing to ESRD, enabling more targeted and timely interventions. This approach aligns with the goals of personalized medicine, where treatment plans are tailored to the specific needs and risk profiles of individual patients. Moreover, the reduced error rates with the updated 2021 eGFR equation suggest an improvement in prediction accuracy for minoritized populations, particularly African American patients, which is a positive step toward reducing healthcare disparities. These findings advocate for continued efforts to refine and optimize predictive models, ensuring equitable and effective care for all patients, regardless of their demographic background.
5.1 Limitations
Our study’s reliance on data from a single institution limits the model’s generalizability to diverse patient populations. Incorporating data from more sources could enhance robustness and applicability across different clinical settings. The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
The use of Electronic Health Records (EHR) data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness. Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions. Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
5.2 Future Directions
Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018. Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial. This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats. This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
Expansion to Other Chronic Diseases: To evaluate our framework’s versatility, we plan to apply it to other chronic conditions, such as diabetes and heart disease. This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
6 Conclusion
As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods. Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression. Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making. The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions. Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management. These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
7 Acknowledgments
We sincerely thank the community nephrology practice for providing access to their de-identified clinical data and the wealth of medical knowledge and inputs shared by several nephrologists in the practice and the CIO on the raw data and the results of our analysis. Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice, partial funding for the study, and the invaluable help provided by their data team in understanding the claims data. We also acknowledge the fellowship support for Y.L from the Center for Machine Learning and Health (CMLH) at Carnegie Mellon University.
8 Conflicts of interest
None declared.

### v3

@@
 Objective: To improve prediction of Chronic Kidney Disease (CKD) progression to End Stage Renal Disease (ESRD) using machine learning (ML) and deep learning (DL) models applied to an integrated clinical and claims dataset of varying observation windows, supported by explainable AI (XAI) to enhance interpretability and reduce bias.
-Materials and Methods: We utilized data about 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018.
+Materials and Methods: We utilized data from 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018.
 Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows.
-Feature importance and Shapley value analysis were employed to understand key predictors.
+Feature importance and SHAP analysis were employed to understand key predictors.
 Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
@@
 A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy.
-The 2021 eGFR equation improved prediction accuracy and reduced racial bias, notably for African American patients.
-Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities.
+The 2021 eGFR equation improved prediction accuracy and reduced racial bias, especially for African American patients.
+Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities for this population.
 Conclusion: This study presents a robust framework for predicting ESRD outcomes in CKD patients, improving clinical decision-making and patient care through multi-sourced, integrated data and AI/ML methods.
@@
 […]
-Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 4 or administrative claims data 5.
+Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 5 or administrative claims data 6; 7.
 While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression.
@@
 For example, Tangri et al.
-7 used clinical variables like age, gender, eGFR, and albuminuria to predict ESRD progression, and Keane et al.
-8 utilized serum creatinine and urine protein levels for high-risk identification.
+9 used clinical variables such as age, gender, estimated glomerular filtration rate (eGFR)10, and albuminuria to predict ESRD progression, and Keane et al.
+11 utilized serum creatinine and urine protein levels for high-risk patient identification.
 Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices.
 Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
-In the field of explainable AI (XAI) for healthcare, established methods like SHapley Additive exPlanations (SHAP) 9 and Local Interpretable Model-agnostic Explanations (LIME) 10 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models.
+In the field of explainable AI (XAI) for healthcare, established methods such as SHapley Additive exPlanations (SHAP) 12 and Local Interpretable Model-agnostic Explanations (LIME) 13 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models.
 However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
@@
 At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus.
-This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making and ultimately improving patient outcomes.
+This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making to ultimately improve patient outcomes.
 2 Objective
-This research aims to evaluate multiple prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data.
-Our primary objective is to estimate the probability that a CKD patient will develop ESRD in the future based on the data from a specified observation window prior to the diagnosis of ESRD, represented as follows:
-$P\left(\mathrm{ESRD}_{i}=1\mid t,X_{\text{clinical},i}(t),X_{\text{claim},i}(t)\right),$
-where $X_{\text{clinical},i}(t)$ and $X_{\text{claim},i}(t)$ denote the clinical and claims data observed up to time $t$ for patient $i$ who progresses to ESRD after the observation window.
-We address two key scenarios in our study:
-  • Model Robustness: We train and test different analytical models on the same dataset to ensure consistency and comparability, allowing us to select the best-performing model with a focus on robustness and reliability.
-  • Clinical Utility: In clinical settings, the likely duration of CKD progression to ESRD is unknown.
-Our model aims to predict the likelihood of ESRD in the future and suggests an optimal observation window based on the duration of data available on the patient from the time of CKD Stage 3 diagnosis in time steps of 6, 12, 18, 24 and 30 months.
-To address these objectives, we focus on identifying the best data-driven observation window while maintaining an acceptable error rate $\epsilon(t)$, represented as follows:
-$T_{\text{obs }}=\min\left\{t\mid\epsilon(t)\leq\epsilon^{\prime},t\in\{6,12,18,24,30\}\right\},\vspace{-.5cm}$
-where $\epsilon^{\prime}$ is defined as:
-$\epsilon^{\prime}=\left\{\begin{array}[]{ll}\textnormal{user-defined error rate}&\textnormal{if specified},\\ \min_{t\in\{6,12,18,24,30\}}\epsilon(t)&\textnormal{otherwise}\end{array}\right.$
+This research aims to evaluate prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data.
+Fig.
+1 illustrates the key temporal components of our study design, with the observation window beginning at Stage 3 diagnosis (t=0).
+During this observation window, patients may progress through different CKD stages (S4, S5); however, we specifically focus on the cohort that does not develop ESRD within the observation window itself.
+This ensures that our predictive modeling uses only pre-ESRD data to forecast future ESRD development.
+Our primary objectives are:
+  • ESRD Risk Prediction: To estimate the probability that a CKD patient will develop ESRD in the future, based on data from a specified observation window that begins from each patient’s first occurrence of CKD stage 3.
+This probability is represented as:
+    $P\left(\mathrm{ESRD}_{i}=1\mid t>t_{\mathrm{obs}},X_{\text{clinical },i}\left(t_{\mathrm{obs}}\right),X_{\text{claim },i}\left(t_{\mathrm{obs}}\right)\right),$
+    where $X_{\text{clinical},i}\left(t_{\text{obs}}\right),X_{\text{claim},i}\left(t_{\text{obs}}\right)$ denote the clinical and claims data observed up to time $t_{\text{obs}}$ for patient $i$.
+  • Observation Window Optimization: To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
+    $T_{\text{obs }}=\arg\max_{t\in\{6,12,18,24,30\}}\text{ Performance }(t),$
+    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score (the harmonic mean of precision and recall) 14, Area Under the Receiver Operating Characteristic curve (AUROC) 15, and Area Under the Precision-Recall Curve (AUPRC) 16.
+Figure 1: A timeline illustration of observation windows for CKD progression to ESRD.
+The observation starts at CKD Stage 3.
+Objective 1 estimates the probability of ESRD occurrence using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
+This approach balances the trade-off between maximizing predictive accuracy and minimizing the duration of the observation window required for reliable predictions.
 Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
-3 Materials and Methods
-3.1 Data
+3 Data Sources
 This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data.
-The de-identified administrative claims dataset, obtained from a major health insurance provider and approved by their Institutional Review Board, includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018.
-It encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with Chronic Kidney Disease (CKD) and managed by a leading community nephrology practice that shared their de-identified clinical data.
-The clinical dataset was sourced from the EHR of the nephrology practice, containing laboratory results, patient demographics, diagnostic details, and medication records from 1998 to 2018.
-For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data.
-Usage of the clinical data was approved by the Institutional Review Board of the University.
-Preprocessing of the claims data involved removing duplicates, excluding records without any stage of CKD diagnoses, and eliminating entries with negative costs to ensure data integrity.
-After cleansing, the claims dataset included 5,317,178 individual claims across 7,129 unique patients.
-The clinical dataset was similarly refined to include 433,421 lab records for 10,326 patients.
-These datasets were then integrated using unique, de-identified patient IDs, providing a comprehensive view of each patient’s medical and claims history.
-This integration was crucial for accurately identifying ESRD predictors using multiple analytical models.
-3.2 Research Framework
+The de-identified administrative claims dataset was provided by Highmark, a major health insurance provider, and includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018.
+This dataset encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD.
+The use of this dataset was approved by Highmark’s Institutional Review Board.
+The clinical dataset was sourced from the EHR data of TMA, P.C., a leading community nephrology practice.
+This dataset contains laboratory results, patient demographics, diagnostic details, and medication records spanning from 1998 to 2018.
+For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data (2009–2018).
+Usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
+Preprocessing of the claims data involved removing duplicates, excluding records without any CKD stage information, and eliminating entries with negative costs to ensure data integrity.
+After cleaning, the claims dataset included 5,317,178 individual claims across 7,129 unique patients.
+The clinical dataset was similarly refined to include 433,421 laboratory records for 10,326 patients.
+These datasets were then integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
+4 Methods
+As shown in Figure 7, our approach to predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) encompasses three primary stages: Data Preparation, Modeling and Validation, and Additional Analyses to assess robustness and clinical utility.
 Figure 2: Overview of research framework and training strategy.
@@
 Part C outlines the modeling strategy.
-As illustrated in Fig.1, our framework for predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) consists of three key stages: Data Preparation, Modeling & Validation, and Additional Analyses form robustness and clinical utility.
-3.2.1 Data Preparation
-Data Pre-processing is the foundation of the framework, involving the integration and thorough cleaning of claims and clinical datasets.
-This Part A stage includes data preprocessing steps such as cleansing, normalization, missing value imputations and log transformations as appropriate for data analysis.
-Cohort Identification: Following this, in Part B, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period.
+4.1 Data Preparation
+4.1.1 Data Pre-processing
+As mentioned earlier, we performed data cleansing to ensure the integrity of the datasets.
+This process involved removing duplicate records to prevent redundancy and potential biases in the analysis.
+We also excluded entries lacking any stage of CKD diagnoses to maintain the datasets’ relevance to our study objectives, and claims records with negative values, which likely indicated data entry errors.
+To address missing data issue, we employed Multiple imputation using chained equation (MICE)17, a widely adopted method in healthcare research for its ability to account for multivariable relationships and provide robust imputations.
+Normalization was applied to ensure that features contributed proportionately to the analysis.
+We scaled numerical values to a standard range, a step essential for algorithms sensitive to the magnitude of input values.
+We applied log transformations 18 to variables exhibiting skewed distributions to stabilize variance and approximate normality.
+This approach is particularly beneficial for handling data with exponential growth patterns, as it mitigates the impact of outliers and enhances the performance of linear models.
+In the laboratory data, we observed an additional challenge where many records lacked unit specifications, leading to significant discrepancies in numerical values.
+To address this, we conducted distribution detection & adjustment19 to identify and correct unit inconsistencies.
+By analyzing the statistical distribution of each laboratory measurement, we detected shifts indicative of differing units.
+In consultation with clinical experts, we established expected value ranges for each measurement.
+Outlier values that deviated markedly from these ranges were adjusted to align with the appropriate units, ensuring consistency and comparability across all records.
+4.1.2 Cohort Identification
+In Part B of Fig.7, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period.
 Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
-The study cohort was curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists.
-We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management.
-Further refinement excluded patients who progressed to ESRD within the observation window, resulting in a final cohort with uninterrupted, detailed medical and claims histories.
-Feature Engineering, depicted in Part C, involved identifying and creating predictive variables essential for modeling.
+The study cohort was further curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists.
+We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management, resulting in a final cohort with uninterrupted, detailed clinical and claims histories.
+4.1.3 Feature Engineering
+Part C depicted in Fig.
+7 involves identifying and creating predictive variables essential for modeling.
 With claims data, we categorized the features used for prediction into two main groups.
@@
 These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
-The second group, comorbidity-based features, includes patient-specific clinical information such as the duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia.
-These features are crucial for understanding CKD progression and were statistically analyzed for their significance in predicting ESRD outcomes (as detailed in Table 1(a) and Table 1(b)).
-Table 1: Summary of Descriptive and Categorical Statistics for CKD Features
-(a) Descriptive statistics for claims-driven and clinical-driven features; means and standard deviations (StD) for ESRD vs.
-non-ESRD populations.
-The $p$-value is based on the t-test of the difference between the ESRD and non-ESRD populations.
-A significance level of $0.05$ is used.
-Category Feature Name Overall Mean (StD) ESRD Mean (StD) Non-ESRD Mean (StD) $p$-value
-Claims-driven Count of pharmacy claims (CM1) 110 (88) 120 (94) 109 (87) 0.035
-Count of inpatient claims (CM2) 3.74 (3.80) 3.75 (3.43) 3.74 (3.87) 0.959
-Count of outpatient claims (CM3) 23.69 (22.07) 27.78 (28.75) 23.09 (20.83) 0.003
-Count of professional claims (CM4) 89.73 (68.53) 105.37 (77.46) 87.43 (68.01) 0.000
-Net cost of pharmacy claims (CM5) 10645 (20303) 12053 (17596) 10440 (20662) 0.110
-Net cost of inpatient claims (CM6) 30038 (36080) 33909 (53540) 29440 (32541) 0.121
-Net cost of outpatient claims (CM7) 8657 (17495) 9354 (17522) 8554 (17492) 0.415
-Net cost of professional claims (CM8) 12137 (13707) 15512 (18657) 11640 (12748) 0.000
-Range of claims costs (CM9) 9173 (15910) 11352 (32606) 8852 (11550) 0.147
-Standard deviation of claims costs (CM10) 766 (879) 831 (1263) 757 (806) 0.277
-Clinical-driven Age at CKD stage 3 first diagnosis (CL1) 72.68 (11) 70.13 (10.37) 73.04 (11) 0.000
-CKD stage 3 duration time (days) (CL2) 660 (179) 543 (250) 677 (160) 0.000
-Number of emergency department visits (CL3) 2.11 (2.94) 2.18 (2.63) 2.01 (2.99) 0.258
-(b) Categorical clinical-driven features; proportions of ESRD vs.
-non-ESRD populations.
-The difference between the ESRD and non-ESRD populations is statistically significant for most of the clinical-driven categorical features based on chi-squared test of independence results with a significance level of $0.05$.
-Category Feature Name ESRD Proportion Non-ESRD Proportion ${\chi}^{2}$ Statistics $p$-value
-Clinical-driven Gender (CL4) (male:female) 60:40 52:48 6.73 0.009
-Occurrence of CKD stage 4 (CL5) 43:57 12:88 217.8 0.000
-Occurrence of CKD stage 5 (CL6) 5:95 1:99 48.8 0.000
-Diabetes (CL7) 73:27 59:41 27.3 0.000
-Anemia (CL8) 64:36 62:38 0.58 0.445
-Metabolic acidosis (CL9) 25:75 18:82 9.41 0.002
-Proteinuria (CL10) 13:87 17:83 3.72 0.054
-Secondary hyperparathyroidism (CL11) 33:67 18:82 42.9 0.000
-Phosphatemia (CL12) 5:95 3:97 3.55 0.06
-Atherosclerosis (CL13) 6:94 14:86 18.04 0.000
-Heart failure (CL14) 5:95 11:89 13.7 0.000
-Stroke (CL15) 1:99 3:97 4.58 0.032
-Conduction & dysrhythmias (CL16) 5:95 16:84 29.9 0.000
-Hypertension (CL17) 99:1 97:3 3.1 0.078
+The second group, comorbidity-based features, includes patient-specific clinical information such as the patient duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia.
 Clinical data features were categorized into three groups based on consultations with nephrologists, commonly referenced literature, and the characteristics of our datasets.
 These groups include demographic features, laboratory data features, and comorbidity features.
-  • Demographic Features: These include age, gender, ethnicity and BMI, which have been widely used for understanding CKD progression.
-  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, bicarbonate, and albumin, which are important for monitoring CKD.
-Notably, UACR was not selected due to insufficient readings compared to the other features.
+  • Demographic Features: These include age, gender, ethnicity and body mass index (BMI), which have been widely used for understanding CKD progression.
+  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate, which are important for monitoring CKD.
+Notably, urine albumin-to-creatinine ratio (UACR) 20 and sodium were not selected due to insufficient readings in the dataset.
   • Additional Comorbidity Features: These capture the presence of additional health conditions such as cardiovascular disease, anemia, and metabolic acidosis, in addition to those identified using claims data, and have been shown to influence CKD outcomes.
-3.2.2 Modeling & Validation
+Following these steps, statistical analyses were performed to compare variables between patients who progressed to ESRD and those who did not.
+For continuous variables, data were assessed for normality; those meeting normality assumptions were analyzed using independent t-tests to compare group means.
+For non-normally distributed variables, logarithmic transformations were applied during preprocessing to approximate normality before conducting t-tests.
+Categorical variables were expressed as frequencies (percentages) and compared using chi-squared tests.
+Statistical significance was set at $p<0.05$ for all analyses.
+4.2 Modeling & Validation
 To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns.
-Our modeling strategy (Fig.
-1 Part C) involved partitioning data into training (60%), validation (20%), and testing (20%) sets to ensure model reliability.
-We optimized performance through hyper-parameter tuning and an iterative process to determine the optimal observation window that balances predictive accuracy and timeliness.
-Machine Learning Methods
-We employed statistical methods, beginning with logistic regression (LR) 11 as a baseline to estimate the probability of ESRD progression using clinical and claims data features.
-Beyond LR, we utilized ML techniques such as Random Forest (RF) 12 and Extreme Gradient Boosting (XGBoost) 13.
+Our overarching strategy involved partitioning the dataset into training, validation, and testing subsets through stratified sampling, maintaining class distribution across all splits.
+This approach is particularly crucial in medical research, where class imbalances can significantly impact model performance.
+To further address the imbalance, we applied SMOTE 21; 22 to the training data, generating synthetic minority samples and enhancing the model’s ability to learn from underrepresented classes.
+4.2.1 Machine Learning Methods
+We employed statistical methods, beginning with logistic regression (LR) 23 as a baseline model to estimate the probability of ESRD progression using clinical and claims data features.
+Beyond LR, we utilized ML techniques such as Random Forest (RF) 24 and Extreme Gradient Boosting (XGBoost) 25.
+We evaluated performance using a standard k-fold cross-validation setup 26.
 RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting.
@@
 This dual approach—cohort-level and individual-level analysis—not only provides insights into the drivers of ESRD progression but also improves model transparency, supporting personalized treatment decisions.
-Deep Learning Methods
-To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 14; 15 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis.
-This approach captures the temporal evolution of CKD, reflecting its gradual progression and aligning with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data.
+4.2.2 Deep Learning Methods
+To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 7; 27 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis.
+These three-month intervals are indexed as timestamps (timestamp 0 for months 0-3, timestamp 1 for months 3-6, and so on), with a 18-month observation window containing six timestamps in total.
+This temporal segmentation approach serves multiple purposes: it captures the temporal evolution of CKD, reflects its gradual progression, and aligns with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data.
 Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression.
-We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 16 for feature extraction, Recurrent Neural Network (RNN) 17, Long Short-Term Memory network (LSTMs 18, Gated Recurrent Unit (GRU) 19 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 20 for sequence modeling.
+This timestamp-based structure enables our model to track the progression of disease markers and identify critical temporal patterns that may indicate increased risk of ESRD development.
+We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 28 for feature extraction, Recurrent Neural Network (RNN) 29, Long Short-Term Memory network (LSTMs 30, Gated Recurrent Unit (GRU) 31 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 32 for sequence modeling.
 These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision.
-Modeling Scenarios: Robustness and Clinical Application
-As outlined in the Objective section, our modeling approach addresses two key scenarios:
-  • Model Robustness: To ensure model consistency and fair comparability, we train and test on a unified dataset.
-We address the issue of shrinking cohort size with longer observation windows by selecting cohorts with a minimum of 30 months of data following CKD stage 3 diagnosis.
-The dataset is partitioned into training (60%), validation (20%), and testing (20%) subsets.
-By varying training data length from 6 to 30 months within this cohort, we assess model performance under different data availability conditions, thereby evaluating robustness and reliability.
-  • Clinical Utility: In clinical practice, the progression timeline from CKD to ESRD is often uncertain.
-To model this uncertainty, we integrate all time-length cohorts, enabling the prediction of ESRD likelihood and optimal observation windows without precise progression timelines.
-This mixed-cohort approach enhances model relevance in real-world clinical settings, accommodating variable patient timelines.
-3.2.3 Additional Analyses
-This stage includes a detailed examination of incorrect predictions, identification of factors contributing to errors, and exploration of potential biases, particularly considering the recent changes to eGFR equations in 2021.
-These analyses are critical for refining the models and ensuring they deliver fair and accurate predictions.
-Analysis of Model Misclassifications
+We forgo k-fold cross-validation due to high computational costs and training variance from random initialization33.
+Modern regularization techniques like dropout reduce this need, so we opted for stratified holdout validation to balance efficiency and reliability while maintaining class distributions.
+4.2.3 Hyperparameter Tuning
+For performance optimization, we employed specialized Python libraries (Python 3.10) to optimize model performance through hyperparameter tuning.
+For machine learning models, we utilized the H2O framework 34, which provides comprehensive grid search capabilities.
+The framework systematically explores a predefined set of hyperparameters to identify the optimal combination for model performance.
+For deep learning models, we leveraged Microsoft’s Neural Network Intelligence (NNI) 35.
+NNI is an open-source Python toolkit that automates neural architecture search and hyperparameter optimization.
+It efficiently dispatches and manages trial jobs generated by various tuning algorithms, facilitating the search for optimal neural architectures and hyperparameter configurations.
+4.3 Additional Analyses
+The comparison of these modeling results helps us identify the optimal observation length and the most reliable and accurate model for predicting CKD progression to ESRD.
+However, recognizing the limitations of focusing solely on fixed timelines, we extended our analysis to address clinical realities.
+In practice, the progression timeline from CKD to ESRD is often uncertain when a patient presents for evaluation.
+To accommodate this variability, we leveraged the selected optimal model and observation window to integrate all time-length cohorts, ranging from 6 to 30 months.
+This integration enables predictions of ESRD likelihood and optimal observation windows without requiring precise progression timelines, enhancing the model’s relevance and applicability in real-world clinical settings.
+Building on this framework, we conducted additional analyses to further refine the model, including a detailed investigation of misclassifications and their potential causes, as well as an evaluation of the impact of updated eGFR equations on racial bias in CKD predictions.
+These efforts aim to ensure fairness and accuracy in the model’s clinical applications.
+4.3.1 Analysis of Model Misclassifications
 To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set.
@@
 By understanding the reasons behind incorrect predictions, we can enhance the model’s ability to support clinicians in making more informed and accurate decisions regarding patient care.
-Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
+4.3.2 Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
 The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias.
-In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare.
-In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 21, which excluded the race coefficient to improve accuracy and equity 22.
-This adjustment aimed to enhance prediction precision, especially for minority populations 23; 24; 21.
-Our analysis, using data from 2009 to 2018, initially applied the older eGFR equation.
+In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of its biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare.
+In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 36, which excluded the race coefficient to improve accuracy and equity 37.
+This adjustment aimed to enhance prediction precision, especially for minority populations 38; 39.
+Our analysis, using data from 2009 to 2018, applied both eGFR equations to the data.
 To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
+It is important to note that adopting the updated equation may alter the cohort used for analysis, as the new eGFR values directly influence CKD stage classification.
+This could impact the cohort identification process, and lead to the reclassification of patients, particularly those near stage boundaries.
+Such changes may affect both the composition of the cohort and the model’s predictions.
 5 Results
-4.1 Data Integration and Cohort for Analysis
-(a) Comparison of CKD stage distribution across claims, clinical, and merged datasets with a 24-month observation window.
-(b) Variation in CKD stage distribution as a function of different observation window lengths.
-Figure 2: Overview of CKD Stage Distribution in Merged Data Across Observation Window Lengths.
-To maximize the utility of multi-sourced integrated data, we merged claims and clinical datasets using unique patient identifiers.
-The distribution of data for individual and merged datasets is presented in Fig.
-2.
-The analysis cohort consists of patients identified at CKD stage 3, marking the beginning of data collection for each patient.
-As shown in Fig.
-2(a), merging the datasets resulted in a significant reduction in the number of patients—from 5,518 and 6,842 in the claims and clinical datasets, respectively, to 1,422 in the merged dataset with a 24-month observation window.
-This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis.
-Fig.
-2(b) illustrates the changes in the total number of patients and their distribution across CKD stages with varying observation window lengths.
+5.1 Cohort for Analysis & Data Characteristics
+(a) Variation in CKD stage distribution as a function of different observation window lengths.
+(b) Comparison of CKD stage distribution across claims, clinical, and merged data with a 30-month observation window.
+Figure 3: Overview of CKD Stage Distribution in Merged Data Across Observation Windows.
+In this subsection, we present the trends observed during our data integration process and provide a detailed overview of the data characteristics for the final cohort utilized in our model performance analysis.
+Fig.3(a) illustrates the changes in patient distribution across CKD stages over varying observation window lengths.
 As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis.
@@
 Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
-4.2 Performance Comparison of Models Across Datasets and Scenarios
-4.2.1 Scenario 1: Model Robustness
-This scenario focuses on evaluating the robustness and reliability of various models by training and testing them on a consistent dataset.
-We varied the training data length from 6 to 30 months within a fixed cohort to assess the models’ performance under different data availability conditions.
-This approach enabled us to evaluate the models’ stability and effectiveness across various observation windows, ensuring consistency and fair comparability.
-Table 2: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 1 (Model Robustness).
-This evaluation focuses on the stability and effectiveness of models across varying observation window lengths.
+To ensure model consistency and fair comparability, we trained and tested on the same cohort.
+Recognizing the challenge of shrinking cohort size with extended observation windows, we focused on patients with at least 30 months of data following CKD stage 3 diagnosis and split the test data from this group.
+For the training data, we used the same cohort but with varying data lengths corresponding to the selected observation windows.
+This approach allowed us to explore the impact of different observation windows and identify the optimal one based on the performance on test data.
+After establishing the cohort with at least 30 months of post-CKD stage 3 data, we compared the datasets derived from claims data only (5,518 patients), clinical data only (6,842 patients), and the merged dataset used for analysis.
+As shown in Fig.
+3(b), merging the datasets resulted in a substantial reduction in the number of patients—to 1,422 patients with a 24-month observation window (which proved optimal in subsequent analyses) and 1,355 patients with a 30-month observation window.
+This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis.
+Table 1: Data characteristics of the patient cohort for analysis
+Characteristics Missing data (%) Progressed to ESRD (n = 86) Non-progressed to ESRD P-value
+(n = 1,336)
+Demographic
+Age (years) 0 69.13 ± 12.37 72.04 ± 11.25 $<$0.001
+Female 0 40 (46.5%) 721 (54.0%) 0.2149
+Race 0 $<$0.001
+White 70 (81.4%) 1242 (93.0%)
+African American 12 (14.0%) 60 (4.5%)
+Others 4 (4.6%) 34 (2.5%)
+BMI 4 28.40 ± 5.32 26.40 ± 6.20 $<$0.001
+Comorbidities
+Hypertension 0 85 (99%) 1,323 (99%) 0.863
+Diabetes 0 63 (73.3%) 788 (59.0%) 0.009
+Anemia 0 55 (64.0%) 828 (62.0%) 0.714
+Metabolic acidosis 0 22 (25.6%) 240 (18.0%) 0.077
+Proteinuria 0 11 (12.8%) 227 (17.0%) 0.312
+Secondary hyperparathyroidism 0 28 (32.6%) 240 (18.0%) $<$0.001
+Phosphatemia 0 4 (4.7%) 40 (3.0%) 0.39
+Heart failure 0 6 (7.0%) 120 (9.0%) 0.526
+Stroke 0 1 (1.2%) 40 (3.0%) 0.506
+Conduction & dysrhythmias 0 4 (4.7%) 214 (16.0%) 0.005
+Claims-driven features
+Count of pharmacy claims 0 120 ± 94 109 ± 86 0.293
+Count of inpatient claims 0 3.85 ± 3.41 3.74 ± 3.62 0.773
+Count of outpatient claims 0 27.78 ± 24.75 22.07 ± 19.13 0.039
+Count of professional claims 0 105.37 ± 77.56 87.43 ± 68.02 0.039
+Net cost of pharmacy claims 0 12053 ± 11596 10440 ± 20662 0.242
+Net cost of inpatient claims 0 33909 ± 53540 29440 ± 32541 0.446
+Net cost of outpatient claims 0 9354 ± 17522 8554 ± 17492 0.682
+Net cost of professional claims 0 15512 ± 18657 11640 ± 12748 0.061
+Range of claims costs 0 11352 ± 32606 8852 ± 11550 0.481
+Standard deviation of claims costs 0 831 ± 1263 757 ± 806 0.593
+Clinical-driven features
+eGFR 0 17.21 ± 5.46 22.78 ± 5.66 $<$0.001
+Hemoglobin 3 12.15 ± 2.19 14.25 ± 1.8 $<$0.001
+Bicarbonate 9 22.9 ± 6.36 25.3 ± 4.22 0.001
+Serum calcium 6 9.39 ± 3.62 10.21 ± 2.86 0.042
+Phosphorus 13 3.61 ± 0.87 3.52 ± 0.72 0.350
+CKD stage 3 duration 0 3.7 ± 0.6 3.9 ± 1.4 0.0009
+Occurrence of CKD stage 4 0 47 (54.7%) 298 (22.3%) $<$0.001
+Occurrence of CKD stage 5 0 42 (48.8%) 277 (20.7%) $<$0.001
+Number of emergency department visits 16 2.63 ± 2.18 2.01 ± 1.99 0.005
+Table 1 summarizes the demographic, comorbidity, claims-driven, and clinical-driven characteristics of the patient cohort, comprising 1,422 patients; 86 progressed to ESRD (6%), 1,336 did not.
+For patients who progressed to ESRD, the mean time from CKD stage 3 to ESRD was 4.82 ± 1.82 years.
+The distribution plot of progression times is provided in Supplementary Figure 1.
+Patients who progressed to ESRD were significantly younger ( 69.13 vs.
+72.04 years, $p<0.001$ ), with notable racial disparities.
+African American patients were more likely to progress to ESRD (14.0% vs.
+4.5%), while White patients had lower progression rates ( $81.4\%$ vs.
+93.0%, $p<0.001$ ).
+Comorbidities such as hypertension were prevalent in both groups (99%), but diabetes was more frequent in the ESRD group ( $73.3\%$ vs.
+$0.00959.0\%,p=$ 0.009 ).
+Secondary hyperparathyroidism and conduction/dysrhythmias were also significantly higher in ESRD patients ( $p<0.001$ and $p=0.005$ , respectively).
+In claims-driven features, ESRD patients showed slightly higher counts of outpatient and professional claims ( $p=0.039$ ) but no significant differences in costs.
+Clinical-driven features highlighted lower eGFR (17.21 vs.
+22.78, $p<0.001$ ), lower hemoglobin levels (12.15 vs.
+$0.00114.25,p<$ 0.001 ), and higher occurrences of advanced CKD stages in the ESRD group.
+These differences underscore the clinical and demographic factors associated with ESRD progression.
+5.2 Performance Comparison of Models Across Datasets
+To facilitate comprehension of model performance across different data sources, we present results using a 24-month observation window, which yielded optimal performance across most models.
+Table 2 compares the predictive performance of traditional machine learning and deep learning models across three scenarios: claims data-only, clinical data-only, and merged data.
+The results demonstrate several key findings across the different data sources.
+In the claims data-only scenario, deep learning models, particularly LSTM (AUROC: 0.92, F1: 0.54) and GRU (AUROC: 0.92, F1: 0.50), substantially outperformed traditional machine learning approaches such as logistic regression (AUROC: 0.72, F1: 0.33) and random forest (AUROC: 0.74, F1: 0.36).
+For clinical data, while the performance gap narrowed, deep learning models maintained their advantage, with LSTM achieving the highest performance (AUROC: 0.88, F1: 0.60).
+Most notably, the merged dataset combining both claims and clinical data yielded the best overall performance, with LSTM achieving better results across all metrics (AUROC: 0.93, F1: 0.65, AUPRC: 0.61).
+Comprehensive results across all models and observation windows are provided in Supplementary Table 1.
+Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data.
+(a) Model performance for prediction of ESRD by using claims data.
+Model Model performance metric
+F1 Score AUROC AUPRC
+Logistic Regression 0.33 0.72 0.44
+Random Forest 0.36 0.74 0.48
+XGBoost 0.39 0.75 0.47
+CNN 0.45 0.82 0.50
+RNN 0.50 0.90 0.52
+LSTM 0.54 0.92 0.55
+GRU 0.50 0.92 0.53
+TCN 0.52 0.88 0.53
+(b) Model performance for prediction of ESRD by using clinical data.
+Model Model performance metric
+F1 Score AUROC AUPRC
+Logistic Regression 0.54 0.76 0.47
+Random Forest 0.58 0.79 0.51
+XGBoost 0.57 0.80 0.52
+CNN 0.56 0.84 0.53
+RNN 0.61 0.85 0.53
+LSTM 0.60 0.88 0.56
+GRU 0.60 0.87 0.55
+TCN 0.61 0.83 0.54
+(c) Model performance for prediction of ESRD by using merged data.
+Model Model performance metric
+F1 Score AUROC AUPRC
+Logistic Regression 0.55 0.75 0.45
+Random Forest 0.60 0.84 0.49
+XGBoost 0.61 0.85 0.51
+CNN 0.56 0.80 0.46
+RNN 0.62 0.87 0.53
+LSTM 0.65 0.93 0.61
+GRU 0.63 0.90 0.58
+TCN 0.61 0.89 0.58
+5.3 Key Feature Analysis Using Explainable AI Techniques
+Having identified the 24-month observation window as optimal and LSTM as the best-performing model for ESRD prediction, we employed two complementary explainable AI techniques to understand the features driving these predictions.
+At the cohort level, we utilized XGBoost’s feature importance analysis to identify key predictive variables across the population.
+At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.
+5.3.1 Cohort-Level Key Feature Identification
+Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.
+4(a)).
+The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD.
+Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly.
+The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power.
+This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice.
+However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
+(a) Top 30 features of importance for the XGBoost model.
+(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
+Figure 4: Key Feature Analysis: Feature importance and SHAP analysis for ESRD risk prediction using a 24-month observation window.
+5.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
+While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach.
+To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
+We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD.
+Despite both having CKD Stage 5, they showed notably different patterns of contributing features, highlighting how the same clinical outcome can arise from distinct combinations of risk factors and underscoring the importance of personalized treatment strategies.
+In Figure 4(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue.
+For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it.
+For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
+This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning.
+SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile and modifiable factors.
+5.4 Analysis of Model Misclassifications: Type I and Type II Errors
+To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 5.
+Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered.
+This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
+To understand the underlying causes of model misclassification, we conducted a detailed analysis of Type I and Type II errors.
+We focused particularly on Type II errors, where patients were incorrectly predicted not to develop ESRD but ultimately did progress to the condition.
+By comparing the 16 false negatives (patients shown in the lower part of the right plot) with the 70 true positives (patients correctly predicted to progress to ESRD, shown in the upper part of the rightplot), we aimed to identify patterns and characteristics that might explain why the model failed to identify these high-risk patients.
+These errors can delay essential treatment, potentially accelerating disease progression.
+Conversely, Type I errors represent patients who were incorrectly predicted to develop ESRD but did not actually progress.
+Our analysis contrasted 17 false positives (patients shown in the upper part of the left plot) with 1,319 true negatives (patients correctly predicted not to progress, shown in the lower part of the left plot).
+These Type I errors are particularly concerning in clinical practice as they could lead to unnecessary medical interventions and inefficient allocation of healthcare resources, potentially diverting attention from patients at genuine risk of ESRD progression.
+Understanding the characteristics distinguishing these misclassified cases from correctly predicted ones may provide valuable insights for model refinement and improved clinical decision support.
+Figure 5: Analysis of Model Misclassification: Type I and Type II Errors.
+The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset.
+The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
+Table 3: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy.
+This table presents representative subsets of features under various timestamps.
+Features in bold appear in both analyses.
+(a) Type II Error Analysis: Feature Comparison Between False Negatives and True Positives.
+Model Mean P-value Timestamp
+Correct Incorrect
+n_claims_I 0.70 0.14 0.048752 5
+n_claims_O 5.71 2.67 0.001325 2
+net_exp_O 3831.83 830.61 0.002747 0
+S5 0.36 0.06 0.001189 6
+S5 0.37 0.49 0.662073 7
+net_exp_O 4219.01 1044.37 0.001883 4
+(b) Type I Error Analysis: Feature Comparison Between False Positives and True Negatives.
+Model Mean P-value Timestamp
+Correct Incorrect
+n_claims_DR 13.52 7.21 0.001156 3
+n_claims_O 3.03 1.69 0.009384 4
+S4 3831.83 830.61 0.002747 4
+S5 0.09 0.35 0.000250 6
+S5 0.08 0.35 0.001388 7
+net_exp_O 1044.82 333.61 0.000324 4
+Based on Table 3(a), CKD Stage 5 emerges as a critical feature.
+For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, which is significantly lower than the 0.36 mean for those correctly predicted to progress.
+This suggests that these patients did not have a Stage 5 record until the subsequent timestamp.
+The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely.
+However, after consulting with nephrologists and examining the literature on CKD progression, we note that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
+Based on Table 3(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD.
+This is unusual.
+One possible explanation could be the issue of data censoring.
+Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later.
+Additional data collection beyond 2018 would be needed to update and refine this analysis.
+5.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
+(a) Incorrect predictive probabilities using the 2009 eGFR equation.
+(b) Incorrect predictive probabilities using the 2021 eGFR equation.
+(c) Demographic distribution patterns under the 2021 eGFR equation. Encoding mappings are: Race: 0=White, 1=Black, 2=Others; Gender: 0=Male, 1=Female; ESRD Status: 0=non-ESRD, 1=ESRD.
+(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
+Figure 6: Impact of Updated eGFR Equations (2009 vs.
+2021) on Prediction Errors and Racial Bias.
+Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
+When comparing predictions between the 2009 and 2021 eGFR equations (Fig  6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228.
+Among these misclassifications, only 7 cases were common to both equations.
+Analysis of racial distribution in prediction errors (Fig  6(c),  6(d)) revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12.
+The demographic breakdown for the 2021 equation cohort, including race, gender, and ESRD outcomes, is shown in Fig  6(c).
+However, these numbers are clearly too small for robust insights.
+6 Discussion
+This study demonstrates the value of integrating multiple data sources and deep learning methodologies for improving ESRD progression predictions in CKD patients.
+By combining claims and EHR data with LSTM and GRU models, which excel at capturing temporal dependencies, we achieved enhanced predictive accuracy using a 24-month observation window.
+This window length provides an optimal balance between early detection and prediction reliability, supporting timely clinical interventions.
+Our research addresses a critical gap in chronic disease management by integrating clinical and claims data, offering a more comprehensive view of patient health trajectories than single-source approaches.
+The combination of medical indicators and healthcare utilization patterns provides deeper insights into ESRD progression risk factors, making the model more robust across diverse patient populations.
+The determination of this optimal observation window represents a significant advancement in balancing timely interventions with prediction accuracy, enabling healthcare providers to efficiently monitor CKD progression while minimizing data collection burden.
+Our findings suggest that while the updated 2021 eGFR equation shows modest improvement in overall prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228) in this study population, the more useful insight may come from the pattern of misclassifications.
+The fact that only 7 cases were misclassified by both equations, despite similar overall error rates, indicates that these equations capture different aspects of kidney function and disease progression.
+This has important implications for clinical practice: while the 2021 equation may better serve minoritized populations by removing race-based adjustments, the persistence of prediction errors clustered in certain patient profiles suggests that eGFR alone may not be sufficient for accurate ESRD prediction.
+These findings underscore the need for a multi-faceted approach to risk assessment, combining both traditional clinical markers and novel predictive features.
+By leveraging integrated data sources and advanced modeling techniques, healthcare providers can develop more comprehensive risk assessment strategies, potentially leading to more targeted interventions and improved patient outcomes across diverse populations.
+6.1 Limitations
+Our study’s reliance on data from a single institution limits the model’s generalizability to all care delivery settings.
+Incorporating data from more sources could enhance robustness and applicability across different clinical settings.
+The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
+The use of EHR data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness.
+Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
+Despite employing data oversampling methods, the disparity between our model’s high AUROC scores and relatively lower F1 and AUPRC scores indicates persistent challenges in addressing data imbalance.
+This gap suggests that traditional oversampling techniques may not fully compensate for the inherent imbalance in ESRD progression data.
+To enhance performance, exploring advanced data balancing methods is essential.
+Techniques such as ensemble methods, cost-sensitive learning, and hybrid approaches that combine oversampling and undersampling have shown promise in effectively managing imbalanced datasets.
+Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions.
+Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
+6.2 Future Directions
+Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018.
+Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial.
+This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
+Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats.
+This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
+Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
+Expansion to Other Chronic Diseases: To evaluate framework’s versatility, we plan to apply it to other chronic conditions, such as heart disease.
+This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
+7 Conclusion
+As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods.
+Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression.
+Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
+A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making.
+The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions.
+Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
+Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management.
+These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
+8 Acknowledgments
+The authors sincerely thank the community nephrology practice for providing access to their de-identified clinical data and the wealth of medical knowledge and inputs shared by several nephrologists in the practice and their CIO on the raw data and the results of our analysis.
+Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice and for the invaluable help provided by their data team in understanding the claims data.
+AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance.
+GPT-4o and Claude 3.5 (Sonnet) were used solely for polishing grammar and refining expression.
+The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity.
+Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
+Funding Statement
+This work was supported by the fellowship support for Y.L. from the Center for Machine Learning and Health (CMLH) at Carnegie Mellon University.
+Competing Interests Statement
+The authors have no competing interests to declare.
+Data Availability Statement
+The data underlying this article consist of active healthcare clinical and claims data, which cannot be shared publicly due to privacy regulations and the sensitive nature of the information.
+Supplementary
+Supplementary Table 1: Model Performance Metrics Across Observation Windows
+Table 4: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data.
 Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
@@
 TCN 0.71 0.49 0.74 0.57 0.88 0.60 0.89 0.61 0.91 0.59
-Table 2(c) compares model performance across claims data-only scenario, clinical data-only, and merged data.
-In claims data-only, ML model XGBoost achieved the highest AUC (0.78) at 18 months, while Random Forest had the best F1 score (0.41).
-DL based LSTM outperformed all others with an AUC of 0.92 and an F1 score of 0.54 at 24 months.
-In clinical data-only scenario, Random Forest led with an AUC of 0.80 and an F1 score of 0.58, but LSTM again excelled with an AUC of 0.88 and an F1 score of 0.63.
-Merged data further improved results, with XGBoost reaching an AUC of 0.85, and LSTM and GRU achieving AUCs of 0.93 and 0.91, with F1 scores of 0.65 and 0.63.
-Overall, merged data and deep learning methods, particularly LSTM and GRU, delivered the best performance, with the best observation window being 24 months.
-4.2.2 Scenario 2: Clinical Application
-In this scenario, we addressed the challenge of predicting ESRD development in real-world clinical settings, where the duration of CKD progression is often uncertain.
-By examining all cohort observation windows, the models predict ESRD likelihood and determine the optimal observation window despite unknown progression timelines.
-This mixed cohort approach enhances the model’s relevance for clinical practice, accommodating variable patient timelines.
-Table 3(c) shows model performance across claims-only, clinical-only, and merged data.
-In claims-only, XGBoost had the highest AUC (0.76) at 12 months among ML models, while Random Forest had the best F1 score (0.40) at 18 months.
-However, LSTM outperformed others with an AUC of 0.90 and an F1 score of 0.51 at 24 months.
-For clinical-only, Random Forest led ML models (AUC 0.79, F1 0.56 at 18 months), but LSTM was superior (AUC 0.85, F1 0.61 at 24 months).
-Merged data yielded the best results, with XGBoost leading ML methods (AUC 0.84 at 18 months), and LSTM and GRU excelling overall (LSTM: AUC 0.90, F1 0.60; GRU: F1 0.63 at 24 months).
-Merged data consistently outperformed individual datasets, and deep learning models, especially LSTM and GRU, showed significant performance gains, highlighting the benefits of integrating multiple data sources for accurate clinical predictions.
-Comparing the two scenarios, we observe that while both benefit from deep learning models and data integration, they demonstrate different strengths depending on the context.
-Scenario 1 underscores the importance of stability and consistency across varying data conditions, whereas Scenario 2 highlights adaptability and practical applicability in real-world clinical settings.
-Together, these scenarios underscore the critical role of deep learning and comprehensive data integration, with 24 months identified as a key observation window that balances robustness and clinical relevance.
-Table 3: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 2 (Clinical Application).
-The model aims to predict the likelihood of ESRD development and suggest an optimal observation window based on data from the patient’s initial stage 3 CKD visit.
-Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
-(a) Performance of models using claims data only.
-6 months 12 months 18 months 24 months 30 months
-Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
-Claims Data Only Modeling Machine Learning Methods Logistic Regression 0.66 0.22 0.67 0.28 0.70 0.32 0.72 0.34 0.71 0.35
-Random Forest 0.72 0.26 0.75 0.30 0.74 0.40 0.73 0.34 0.68 0.31
-XGBoost 0.72 0.28 0.76 0.31 0.72 0.36 0.71 0.32 0.69 0.33
-Deep Learning Methods CNN 0.70 0.34 0.72 0.48 0.79 0.49 0.88 0.46 0.86 0.42
-RNN 0.79 0.43 0.78 0.47 0.82 0.47 0.84 0.49 0.82 0.41
-LSTM 0.79 0.33 0.88 0.33 0.84 0.47 0.90 0.51 0.87 0.41
-GRU 0.73 0.29 0.85 0.44 0.86 0.50 0.89 0.49 0.86 0.40
-TCN 0.72 0.42 0.82 0.38 0.84 0.48 0.87 0.49 0.85 0.40
-(b) Performance of models using clinical data only.
-6 months 12 months 18 months 24 months 30 months
-Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
-Clinical Data Only Modeling Machine Learning Methods Logistic Regression 0.61 0.45 0.69 0.48 0.78 0.49 0.76 0.50 0.73 0.51
-Random Forest 0.72 0.51 0.75 0.53 0.79 0.56 0.73 0.54 0.68 0.50
-XGBoost 0.72 0.51 0.76 0.54 0.78 0.58 0.71 0.56 0.69 0.57
-Deep Learning Methods CNN 0.70 0.44 0.72 0.46 0.77 0.46 0.80 0.56 0.76 0.54
-RNN 0.74 0.48 0.72 0.44 0.78 0.48 0.82 0.59 0.78 0.54
-LSTM 0.76 0.44 0.78 0.46 0.80 0.51 0.85 0.61 0.79 0.55
-GRU 0.73 0.46 0.78 0.52 0.80 0.50 0.85 0.59 0.77 0.47
-TCN 0.69 0.47 0.75 0.48 0.77 0.48 0.81 0.59 0.77 0.51
-(c) Performance of models using merged data.
-6 months 12 months 18 months 24 months 30 months
-Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
-Merged Data Modeling Machine Learning Methods Logistic Regression 0.66 0.44 0.68 0.47 0.74 0.51 0.72 0.51 0.65 0.48
-Random Forest 0.72 0.47 0.76 0.51 0.80 0.56 0.81 0.52 0.77 0.53
-XGBoost 0.71 0.50 0.79 0.52 0.84 0.57 0.81 0.56 0.79 0.57
-Deep Learning Methods CNN 0.60 0.43 0.66 0.43 0.77 0.47 0.78 0.48 0.76 0.50
-RNN 0.66 0.49 0.69 0.44 0.80 0.50 0.84 0.55 0.80 0.52
-LSTM 0.68 0.49 0.75 0.48 0.89 0.57 0.90 0.60 0.79 0.53
-GRU 0.65 0.49 0.76 0.49 0.89 0.55 0.90 0.63 0.79 0.52
-TCN 0.65 0.45 0.74 0.50 0.78 0.44 0.88 0.59 0.78 0.53
-4.3 Key Factor Analysis Using Explainable AI Techniques
-We identified a 24-month observation window as optimal for balancing early detection and prediction accuracy, with LSTM proving most effective among deep learning models and XGBoost leading in machine learning.
-Focusing on Scenario 2 for its relevance to real-world clinical settings with uncertain CKD progression timelines, we refined our approach further to enhance clinical applicability.
-We used cohort-level feature importance analysis with XGBoost to determine key predictive features and applied SHAP analysis at the individual patient level to improve model interpretability and transparency.
-4.3.1 Cohort-Level Key Feature Identification
-Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.
-3(a)).
-The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its significance in predicting ESRD.
-Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly.
-The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power.
-This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice.
-However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
-(a) Top 30 features of importance for the XGBoost model.
-(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
-Figure 3: Key Feature Analysis: Feature Importance and SHAP Analysis for ESRD Risk Prediction Using A 24-month Observation Window.
-4.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
-While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach.
-To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
-We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD.
-Despite both having CKD Stage 5, the contributing features differed significantly, underscoring the need for personalized treatment strategies.
-In Figure 3(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue.
-For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it.
-For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
-This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning.
-SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile, thereby enhancing the effectiveness of clinical care.
-4.4 Analysis of Model Misclassifications: Type I and Type II Errors
-To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 4(a).
-Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered.
-This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
-To understand what drives these misclassifications, we conducted a detailed analysis of Type I and Type II errors, illustrated in Figures 4(c) and 4(b).
-Type II errors (Figure 4(b)) involve patients wrongly predicted not to develop ESRD but did; we compared these 16 false negatives to the 70 correctly predicted to progress to ESRD.
-These errors can delay essential treatment, potentially accelerating disease progression.
-Conversely, Type I errors (Figure 4(c)) involve patients wrongly predicted to develop ESRD who did not.
-We contrasted 17 false positives with 1,319 correct predictions of non-progression, highlighting how these errors can lead to unnecessary medical interventions, potentially depriving resources from those who truly need them.
-By examining the feature distributions between these correctly and incorrectly predicted groups, we aimed to identify statistically significant differences that could explain the model’s errors, providing insights for refining the model and improving its predictive accuracy.
-(a) Distribution of Predicted Probabilities for Patient Outcome.
-(b) Analysis of Type II Errors: Comparison of Patients Predicted Not to Develop ESRD Who Actually Did.
-(c) Analysis of Type I Errors: Comparison of Patients Predicted to Develop ESRD Who Actually Did Not.
-Figure 4: Analysis of Model Misclassifications: Type I and Type II Errors.
-The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset.
-The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
-Table 4: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy.
-This table presents representative subsets of features under various timestamps.
-Features in bold appear in both analyses.
-(a) Type II Error Analysis: Feature Comparison Between False Negatives and True Positives.
-Model Mean P-value Significant Timestamp
-Correct incorrect
-n_claims_I 0.70 0.14 0.048752 Yes 5
-n_claims_O 5.71 2.67 0.001325 Yes 2
-net_exp_O 3831.83 830.61 0.002747 Yes 0
-S5 0.36 0.06 0.001189 Yes 6
-S5 0.37 0.49 0.662073 No 7
-net_exp_O 4219.01 1044.37 0.001883 Yes 4
-(b) Type I Error Analysis: Feature Comparison Between False Positives and True Negatives.
-Model Mean P-value Significant Timestamp
-Correct incorrect
-n_claims_DR 13.52 7.21 0.001156 Yes 3
-n_claims_O 3.03 1.69 0.009384 Yes 4
-S4 3831.83 830.61 0.002747 Yes 4
-S5 0.09 0.35 4.059934e-02 Yes 6
-S5 0.08 0.35 3.911497e-02 Yes 7
-net_exp_O 1044.82 333.61 0.000324 Yes 4
-Based on Table 4(a), CKD Stage 5 emerges as a critical feature.
-For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, significantly lower than the 0.36 mean for those correctly predicted to progress.
-This suggests that these patients did not have a Stage 5 record until the subsequent timestamp.
-The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely.
-However, after consulting with nephrologists and examining the literature on CKD progression to ESRD, we understand that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
-Based on Table 4(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD.
-This is unusual.
-One possible explanation could be the issue of data censoring.
-Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later.
-Additional data collection beyond 2018 would be needed to update and refine this analysis.
-4.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
-(a) Incorrect predictive probabilities using the 2009 eGFR equation.
-(b) Incorrect predictive probabilities using the 2021 eGFR equation.
-(c) Distribution of race, gender, and ESRD status with the 2021 eGFR equation. Categories: Race (0 = White, 1 = Black, 2 = Other), Gender (0 = Male, 1 = Female), ESRD (0 = No ESRD, 1 = ESRD).
-(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
-Figure 5: Impact of Updated eGFR Equations (2009 vs.
-2021) on Prediction Errors and Racial Bias.
-Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
-When comparing the incorrect predictions using the 2009 eGFR equation (Fig  5(a)) with those from the 2021 equation (Fig  5(b)), we observed a decrease in the number of incorrect predictions from 33 to 28.
-This reduction reflects a slight improvement in the error rate, which decreased from 0.0232 to 0.0228 with the updated 2021 eGFR equation.
-Despite this improvement, the pattern of incorrect predictions being clustered towards the lower end, as discussed earlier, persists.
-To assess whether the updated eGFR equation improves prediction accuracy for minoritized groups, we analyzed the demographic and racial distribution of patients who were incorrectly predicted not to develop ESRD.
-Fig  5(c) displays the demographic breakdown, including race, gender, and ESRD outcomes, for the newly defined cohort under the 2021 equation.
-Fig  5(d) reveals changes in incorrect predictions: the number of misclassified white patients increased from 2 to 4, while incorrect predictions for African American patients decreased significantly from 4 to 1.
-Meanwhile, the count for patients in the ”Other” category rose from 10 to 12.
-Given the relatively small sample size for the ”Other” racial category, the increase may not be statistically significant.
-However, the notable reduction in incorrect predictions for African American patients suggests that the 2021 eGFR equation enhances the model’s prediction accuracy for this group.
-This improvement is particularly important as it indicates progress towards reducing racial bias in ESRD predictions, aligning with broader efforts to enhance equity in healthcare.
-Overall, while the updated 2021 eGFR equation has led to a modest reduction in the overall error rate and appears to improve prediction accuracy for African American patients, further analysis with larger cohorts are needed to assess its impact on other racial groups.
-This finding underscores the ongoing need to refine prediction models to ensure fair and accurate outcomes across all patient demographics.
-5 Discussion
-This study underscores the value of integrating patient data from multiple sources, such as claims and EHRs, and of different types, such as clinical comorbidities and utilization costs, combined with deep learning methodologies, particularly in the longitudinal representation of healthcare data, for improving the accuracy of patient outcome predictions.
-By employing models such as LSTM and GRU, which are adept at capturing temporal dependencies in sequential data, we were able to enhance the predictive power to assess progression to ESRD in CKD patients.
-The use of a 24-month observation window, identified as the optimal period for accurate predictions, allows for timely and reliable risk assessment, providing a critical balance between early detection and maintaining a manageable error rate.
-These findings add to the rapidly growing evidence for driving more precise and interpretable predictions for chronic disease management, ultimately supporting better clinical decision-making to improve patient outcomes.
-This research bridges a significant gap in risk assessment for chronic disease management studies by integrating clinical and claims data.
-Previous models have often relied on one data type in isolation, either clinical or claims, which can lead to incomplete representations of a patient’s health journey and the progression of CKD.
-By combining both data types, our approach offers a more comprehensive view, capturing not only medical indicators but also healthcare utilization patterns, as demonstrated by the features that are identified as most important in ESRD prediction.
-This integrated data strategy enhances the model’s robustness and applicability across diverse patient populations, addressing socio-economic factors and healthcare interactions that single-source models may overlook.
-Our findings highlight the importance of a multi-faceted approach to data integration, which can lead to more inclusive and accurate models for managing chronic diseases such as CKD.
-A key innovation of our work is the determination of an optimal observation window that minimizes the volume of data and time required for making reliable predictions while maintaining an acceptable error rate.
-This focus on observation period optimization is crucial for balancing the need for timely interventions with the risks of premature or inaccurate predictions.
-By rigorously evaluating different observation window lengths, we were able to identify a 24-month period as optimal, providing a critical framework for clinicians to monitor and manage CKD progression effectively.
-This approach supports a proactive stance in patient care, allowing for earlier detection and management of patients at risk of progressing to ESRD, thereby potentially reducing long-term healthcare costs and improving patient outcomes.
-The insights derived from our research have significant practical implications for early intervention strategies in CKD management.
-By leveraging integrated data and advanced modeling techniques, healthcare providers can better identify patients at high risk of advancing to ESRD, enabling more targeted and timely interventions.
-This approach aligns with the goals of personalized medicine, where treatment plans are tailored to the specific needs and risk profiles of individual patients.
-Moreover, the reduced error rates with the updated 2021 eGFR equation suggest an improvement in prediction accuracy for minoritized populations, particularly African American patients, which is a positive step toward reducing healthcare disparities.
-These findings advocate for continued efforts to refine and optimize predictive models, ensuring equitable and effective care for all patients, regardless of their demographic background.
-5.1 Limitations
-Our study’s reliance on data from a single institution limits the model’s generalizability to diverse patient populations.
-Incorporating data from more sources could enhance robustness and applicability across different clinical settings.
-The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
-The use of Electronic Health Records (EHR) data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness.
-Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
-Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions.
-Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
-5.2 Future Directions
-Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018.
-Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial.
-This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
-Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats.
-This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
-Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
-Expansion to Other Chronic Diseases: To evaluate our framework’s versatility, we plan to apply it to other chronic conditions, such as diabetes and heart disease.
-This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
-6 Conclusion
-As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods.
-Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression.
-Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
-A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making.
-The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions.
-Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
-Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management.
-These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
-7 Acknowledgments
-We sincerely thank the community nephrology practice for providing access to their de-identified clinical data and the wealth of medical knowledge and inputs shared by several nephrologists in the practice and the CIO on the raw data and the results of our analysis.
-Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice, partial funding for the study, and the invaluable help provided by their data team in understanding the claims data.
-We also acknowledge the fellowship support for Y.L from the Center for Machine Learning and Health (CMLH) at Carnegie Mellon University.
-8 Conflicts of interest
-None declared.
+Supplementary Figure 1
+Figure 7: Distribution of time to ESRD for the ESRD cohort.
+The plot illustrates the variability in time to ESRD across the cohort, with a mean time of 4.82 years and a standard deviation of 1.82 years.
+Supplementary Table 2
+Table 5: Mapping between Feature Abbreviations in The Feature Importance Plot and Full Names
+Short Name Full Name
+S5 Serum Albumin
+MA Metabolic Acidosis
+Diabetes Diabetes
+n_claims_O_min Number of Outpatient Claims (Min)
+Gender Gender
+net_exp_O_std Net Expense for Outpatient Claims (Std Dev)
+n_claims_I_std Number of Inpatient Claims (Std Dev)
+MI Myocardial Infarction
+net_exp_I_max Net Expense for Inpatient Claims (Max)
+Athsc Atherosclerosis
+SH Secondary Hyperparathyroidism
+Phos Phosphatemia
+net_exp_O_min Net Expense for Outpatient Claims (Min)
+ND Nutritional Disorders
+net_exp_DR_max Net Expense for Drug-Related Claims (Max)
+Hemoglobin_max Hemoglobin Level (Max)
+Serum_Calcium_max Serum Calcium Level (Max)
+Phosphorus_std Phosphorus Level (Std Dev)
+Anemia Anemia Diagnosis
+Cvd Cardiovascular Disease
+Intact_PTH_std Intact Parathyroid Hormone (Std Dev)
+Prot Proteinuria
+Age Age
+Phosphorus_max Phosphorus Level (Max)
+n_claims_DR_min Number of Drug-Related Claims (Min)
+net_exp_P_min Net Expense for Pharmacy Claims (Min)
+n_claims_P_std Number of Pharmacy Claims (Std Dev)
+CD Conduction & Dysrhythmias
+Serum_Calcium_min Serum Calcium Level (Min)
+net_exp_I_std Net Expense for Inpatient Claims (Std Dev)

### v4

@@
 Abstract
-Objective: To improve prediction of Chronic Kidney Disease (CKD) progression to End Stage Renal Disease (ESRD) using machine learning (ML) and deep learning (DL) models applied to an integrated clinical and claims dataset of varying observation windows, supported by explainable AI (XAI) to enhance interpretability and reduce bias.
-Materials and Methods: We utilized data from 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018.
-Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows.
+Objective: To improve prediction of Chronic Kidney Disease progression to End Stage Renal Disease using machine learning and deep learning models applied to integrated clinical and claims data with varying observation windows, supported by explainable AI to enhance interpretability and reduce bias.
+Materials and Methods: We utilized data from 10,326 CKD patients, combining clinical and claims information from 2009-2018.
+After preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using five distinct observation windows.
 Feature importance and SHAP analysis were employed to understand key predictors.
-Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
-Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65).
-A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy.
-The 2021 eGFR equation improved prediction accuracy and reduced racial bias, especially for African American patients.
-Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities for this population.
-Conclusion: This study presents a robust framework for predicting ESRD outcomes in CKD patients, improving clinical decision-making and patient care through multi-sourced, integrated data and AI/ML methods.
-Future research will expand data integration and explore the application of this framework to other chronic diseases.
+Models were tested for robustness, clinical relevance, misclassification patterns, and bias.
+Results: Integrated data models outperformed single data source models, with Long Short-Term Memory (LSTM) achieving the highest AUC (0.93) and F1 score (0.65).
+A 24-month observation window optimally balanced early detection and prediction accuracy.
+The 2021 eGFR equation improved prediction accuracy and reduced racial bias, particularly for African American patients.
+Discussion: Improved prediction accuracy, interpretability and bias mitigation strategies have the potential to enhance CKD management, support targeted interventions, and reduce healthcare disparities.
+Conclusion: This study presents a robust framework for predicting ESRD outcomes, improving clinical decision-making through integrated multi-sourced data and advanced analytics.
+Future research will expand data integration and extend this framework to other chronic diseases.
 1 Introduction
 […]
-Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 5 or administrative claims data 6; 7.
-While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression.
-For instance, Sharma et al.
-8 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al.
-6 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database.
-However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics.
-For example, Tangri et al.
-9 used clinical variables such as age, gender, estimated glomerular filtration rate (eGFR)10, and albuminuria to predict ESRD progression, and Keane et al.
+Previous CKD progression prediction efforts used either EHR clinical data 5 or administrative claims 6; 7.
+These approaches often use limited features, not fully capturing CKD progression complexity.
+Sharma et al.
+8 developed a claims-based model to identify CKD patients at Hyperkalemia risk, while Krishnamurthy et al.
+6 predicted CKD onset using comorbidities and medications.
+Claims data typically lacks clinical data’s granularity.
+Tangri et al.
+9 used age, gender, and eGFR 10 to predict ESRD progression, and Keane et al.
 11 utilized serum creatinine and urine protein levels for high-risk patient identification.
-Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices.
-Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
-In the field of explainable AI (XAI) for healthcare, established methods such as SHapley Additive exPlanations (SHAP) 12 and Local Interpretable Model-agnostic Explanations (LIME) 13 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models.
-However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
-This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets.
-At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus.
-This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making to ultimately improve patient outcomes.
+Models using only clinical data may miss complete healthcare system interactions, struggle with missing data and inconsistent recording, and overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse populations.
+This study bridges a critical gap by developing a framework that utilizes integrated clinical and claims data rather than isolated data sources.
+By minimizing the observation window needed for accurate predictions, our approach balances clinical relevance with patient-centered practicality.
+This integration enhances both predictive accuracy and clinical utility, enabling more informed decision-making to improve patient outcomes.
 2 Objective
-This research aims to evaluate prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data.
+This research evaluates predictive models for CKD progression to ESRD using integrated clinical and claims data.
 Fig.
-1 illustrates the key temporal components of our study design, with the observation window beginning at Stage 3 diagnosis (t=0).
-During this observation window, patients may progress through different CKD stages (S4, S5); however, we specifically focus on the cohort that does not develop ESRD within the observation window itself.
-This ensures that our predictive modeling uses only pre-ESRD data to forecast future ESRD development.
-Our primary objectives are:
-  • ESRD Risk Prediction: To estimate the probability that a CKD patient will develop ESRD in the future, based on data from a specified observation window that begins from each patient’s first occurrence of CKD stage 3.
-This probability is represented as:
+1 illustrates our study design, where the observation window begins at the initial diagnosis of CKD Stage 3 (t=0).
+Although patients may progress through subsequent CKD stages (Stages 4 and 5) during this observation period, our cohort specifically excludes those developing ESRD within this timeframe.
+This ensures that predictive modeling utilizes only pre-ESRD data for forecasting future ESRD onset.
+  • Our primary objective, ESRD risk prediction, aims to estimate the probability that a patient diagnosed with CKD Stage 3 will develop ESRD after the observation window, represented as follows:
     $P\left(\mathrm{ESRD}_{i}=1\mid t>t_{\mathrm{obs}},X_{\text{clinical },i}\left(t_{\mathrm{obs}}\right),X_{\text{claim },i}\left(t_{\mathrm{obs}}\right)\right),$
     where $X_{\text{clinical},i}\left(t_{\text{obs}}\right),X_{\text{claim},i}\left(t_{\text{obs}}\right)$ denote the clinical and claims data observed up to time $t_{\text{obs}}$ for patient $i$.
-  • Observation Window Optimization: To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
+  • To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
     $T_{\text{obs }}=\arg\max_{t\in\{6,12,18,24,30\}}\text{ Performance }(t),$
-    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score (the harmonic mean of precision and recall) 14, Area Under the Receiver Operating Characteristic curve (AUROC) 15, and Area Under the Precision-Recall Curve (AUPRC) 16.
+    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score 12, Area Under the Receiver Operating Characteristic curve (AUROC) 13, and Area Under the Precision-Recall Curve (AUPRC) 14.
 Figure 1: A timeline illustration of observation windows for CKD progression to ESRD.
 The observation starts at CKD Stage 3.
-Objective 1 estimates the probability of ESRD occurrence using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
-This approach balances the trade-off between maximizing predictive accuracy and minimizing the duration of the observation window required for reliable predictions.
-Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
+Objective 1 estimates the probability of ESRD occurrence after the observation window, using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
 3 Data Sources
-This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data.
-The de-identified administrative claims dataset was provided by Highmark, a major health insurance provider, and includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018.
-This dataset encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD.
-The use of this dataset was approved by Highmark’s Institutional Review Board.
-The clinical dataset was sourced from the EHR data of TMA, P.C., a leading community nephrology practice.
-This dataset contains laboratory results, patient demographics, diagnostic details, and medication records spanning from 1998 to 2018.
-For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data (2009–2018).
-Usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
-Preprocessing of the claims data involved removing duplicates, excluding records without any CKD stage information, and eliminating entries with negative costs to ensure data integrity.
-After cleaning, the claims dataset included 5,317,178 individual claims across 7,129 unique patients.
-The clinical dataset was similarly refined to include 433,421 laboratory records for 10,326 patients.
-These datasets were then integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
+This study utilized two integrated datasets: administrative claims data and clinical data.
+The claims dataset includes patient healthcare interactions, spanning from 2009 to 2018, containing diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD.
+The clinical dataset from electronic health records contains laboratory results, patient demographics, diagnostic details, and medication records, which was truncated to match the 10-year span of the claims data for consistency of integration.
+See Appendix A for data details.
 4 Methods
-As shown in Figure 7, our approach to predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) encompasses three primary stages: Data Preparation, Modeling and Validation, and Additional Analyses to assess robustness and clinical utility.
-Figure 2: Overview of research framework and training strategy.
-Part A provides an overview of the data preparation, modeling and validation, and additional analyses pipeline.
-Part B illustrates the cohort identification process, detailing the inclusion and exclusion criteria applied to claims and clinical records.
-Part C outlines the modeling strategy.
+As illustrated in Figure 2(a), our comprehensive methodology for predicting CKD progression to ESRD consists of three primary stages: Data Preparation, Modeling, and Additional Analyses.
+Within the Modeling stage, we employ a two-phase analytical framework that first systematically evaluates various predictive models across multiple observation windows (but the same patient cohort for consistent comparisons) to determine optimal performance, and subsequently retrains the most effective model on the cohort for the each observation window for practical clinical application.
+This framework bridges statistical rigor and clinical relevance, facilitating actionable insights for nephrologists managing patients with CKD stage 3 and above.
+(a) Three-stage pipeline for predicting CKD progression to ESRD.
+Stage 1: feature engineering, data preprocessing, and cohort identification using clinical and claims data.
+Stage 2: (2A) systematic evaluation of observation windows with ML/DL models; (2B) retraining the optimal model on the filtered cohort from the optimal observation window, followed by explainable AI analysis.
+Stage 3: Additional analyses including bias detection and misclassification assessment.
+(b) Multi-step cohort curation process for longitudinal CKD analysis.
+Starting with initial selection of patients at CKD stage 3, an adjustable observation window (18 months shown) was applied.
+Only patients with complete clinical and claims data throughout this period were retained, while those progressing to ESRD within the observation window were excluded, ensuring a cohort for analysis (highlighted in yellow) with uninterrupted longitudinal records.
+Figure 2: Framework for CKD progression prediction: (A) Overview of the three-stage analytical pipeline, and (B) Cohort identification methodology.
 4.1 Data Preparation
 4.1.1 Data Pre-processing
-As mentioned earlier, we performed data cleansing to ensure the integrity of the datasets.
-This process involved removing duplicate records to prevent redundancy and potential biases in the analysis.
-We also excluded entries lacking any stage of CKD diagnoses to maintain the datasets’ relevance to our study objectives, and claims records with negative values, which likely indicated data entry errors.
-To address missing data issue, we employed Multiple imputation using chained equation (MICE)17, a widely adopted method in healthcare research for its ability to account for multivariable relationships and provide robust imputations.
-Normalization was applied to ensure that features contributed proportionately to the analysis.
-We scaled numerical values to a standard range, a step essential for algorithms sensitive to the magnitude of input values.
-We applied log transformations 18 to variables exhibiting skewed distributions to stabilize variance and approximate normality.
-This approach is particularly beneficial for handling data with exponential growth patterns, as it mitigates the impact of outliers and enhances the performance of linear models.
-In the laboratory data, we observed an additional challenge where many records lacked unit specifications, leading to significant discrepancies in numerical values.
-To address this, we conducted distribution detection & adjustment19 to identify and correct unit inconsistencies.
-By analyzing the statistical distribution of each laboratory measurement, we detected shifts indicative of differing units.
-In consultation with clinical experts, we established expected value ranges for each measurement.
-Outlier values that deviated markedly from these ranges were adjusted to align with the appropriate units, ensuring consistency and comparability across all records.
+We removed duplicate records to prevent redundancy and bias in the analysis.
+Entries lacking CKD diagnoses were excluded to maintain relevance to study objectives, along with claims containing negative values (likely data entry errors).
+The cleaned claims dataset comprised 5,317,178 claims across 7,129 unique patients, while the refined clinical dataset included 433,421 laboratory records for 10,326 patients.
+These datasets were integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
+To address missing data, we employed multiple imputation via chained equations (MICE) 15.
+Continuous numerical features were standardized, and skewed variables were log-transformed 16 to stabilize variance and mitigate outliers.
+For laboratory values missing unit specifications, we used distribution detection and adjustment 17 to identify outliers and harmonize units across records (medical expert-advised ranges in Appendix B).
 4.1.2 Cohort Identification
-In Part B of Fig.7, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period.
-Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
-The study cohort was further curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists.
-We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management, resulting in a final cohort with uninterrupted, detailed clinical and claims histories.
+As depicted in Fig.
+2(b), cohorts were curated through a multi‐step process: patients at CKD stage 3 were selected and an observation window—18 months in the example, but adjustable per study—was applied.
+Only individuals with complete clinical and claims data throughout this window were retained, and those who progressed to ESRD within it were excluded, yielding a cohort for analysis.
 4.1.3 Feature Engineering
-Part C depicted in Fig.
-7 involves identifying and creating predictive variables essential for modeling.
-With claims data, we categorized the features used for prediction into two main groups.
-The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision.
-These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
-The second group, comorbidity-based features, includes patient-specific clinical information such as the patient duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia.
-Clinical data features were categorized into three groups based on consultations with nephrologists, commonly referenced literature, and the characteristics of our datasets.
-These groups include demographic features, laboratory data features, and comorbidity features.
-  • Demographic Features: These include age, gender, ethnicity and body mass index (BMI), which have been widely used for understanding CKD progression.
-  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate, which are important for monitoring CKD.
-Notably, urine albumin-to-creatinine ratio (UACR) 20 and sodium were not selected due to insufficient readings in the dataset.
-  • Additional Comorbidity Features: These capture the presence of additional health conditions such as cardiovascular disease, anemia, and metabolic acidosis, in addition to those identified using claims data, and have been shown to influence CKD outcomes.
-Following these steps, statistical analyses were performed to compare variables between patients who progressed to ESRD and those who did not.
-For continuous variables, data were assessed for normality; those meeting normality assumptions were analyzed using independent t-tests to compare group means.
-For non-normally distributed variables, logarithmic transformations were applied during preprocessing to approximate normality before conducting t-tests.
+Feature engineering involved identifying predictive variables essential for modeling.
+For claims data, we derived two main groups of features: cost-based features, including claim counts, aggregate patient costs, cost ranges, and cost standard deviations from inpatient, outpatient, professional, pharmacy, and vision claims; and comorbidity-based features, such as CKD stage 3 duration, emergency department visit frequency, and critical comorbidities (e.g., hypertension, diabetes, phosphatemia).
+Clinical data features were categorized based on nephrology expert consultations, relevant literature, and dataset characteristics into demographic features, including age, gender, ethnicity, and BMI as established CKD progression indicators; laboratory features comprising essential clinical parameters like eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate (excluding UACR 18 and sodium due to insufficient data availability); and additional comorbidity features, capturing conditions such as cardiovascular disease, anemia, and metabolic acidosis, which are known to influence CKD outcomes.
+Statistical analyses compared variables between patients progressing to ESRD and those who did not.
+For continuous variables, data were assessed for normality; normally distributed variables were analyzed using independent t-tests to compare group means.
+Non-normally distributed variables underwent logarithmic transformation during preprocessing before applying t-test.
 Categorical variables were expressed as frequencies (percentages) and compared using chi-squared tests.
@@
 4.2 Modeling & Validation
-To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns.
-Our overarching strategy involved partitioning the dataset into training, validation, and testing subsets through stratified sampling, maintaining class distribution across all splits.
-This approach is particularly crucial in medical research, where class imbalances can significantly impact model performance.
-To further address the imbalance, we applied SMOTE 21; 22 to the training data, generating synthetic minority samples and enhancing the model’s ability to learn from underrepresented classes.
+We trained both ML and DL models on stratified train/validation/test splits to preserve class proportions, and applied SMOTE 19; 20 to the training data to oversample the minority class.
 4.2.1 Machine Learning Methods
-We employed statistical methods, beginning with logistic regression (LR) 23 as a baseline model to estimate the probability of ESRD progression using clinical and claims data features.
-Beyond LR, we utilized ML techniques such as Random Forest (RF) 24 and Extreme Gradient Boosting (XGBoost) 25.
-We evaluated performance using a standard k-fold cross-validation setup 26.
-RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting.
-XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
-To enhance model interpretability, we conducted cohort-level feature importance analysis to identify key predictors across the dataset.
-Additionally, SHAP analysis was applied at the individual level to elucidate each feature’s contribution to specific predictions.
-This dual approach—cohort-level and individual-level analysis—not only provides insights into the drivers of ESRD progression but also improves model transparency, supporting personalized treatment decisions.
+We employed logistic regression (LR) 21 as our baseline model for ESRD progression probability estimation.
+We extended our analysis to Random Forest (RF) 22 and XGBoost 23, evaluating performance via k-fold cross-validation 24.
+RF utilizes multiple decision trees to enhance accuracy while reducing overfitting.
+XGBoost sequentially builds models to correct previous errors, proving highly effective for structured data.
 4.2.2 Deep Learning Methods
-To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 7; 27 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis.
-These three-month intervals are indexed as timestamps (timestamp 0 for months 0-3, timestamp 1 for months 3-6, and so on), with a 18-month observation window containing six timestamps in total.
-This temporal segmentation approach serves multiple purposes: it captures the temporal evolution of CKD, reflects its gradual progression, and aligns with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data.
-Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression.
-This timestamp-based structure enables our model to track the progression of disease markers and identify critical temporal patterns that may indicate increased risk of ESRD development.
-We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 28 for feature extraction, Recurrent Neural Network (RNN) 29, Long Short-Term Memory network (LSTMs 30, Gated Recurrent Unit (GRU) 31 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 32 for sequence modeling.
-These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision.
-We forgo k-fold cross-validation due to high computational costs and training variance from random initialization33.
-Modern regularization techniques like dropout reduce this need, so we opted for stratified holdout validation to balance efficiency and reliability while maintaining class distributions.
-4.2.3 Hyperparameter Tuning
-For performance optimization, we employed specialized Python libraries (Python 3.10) to optimize model performance through hyperparameter tuning.
-For machine learning models, we utilized the H2O framework 34, which provides comprehensive grid search capabilities.
-The framework systematically explores a predefined set of hyperparameters to identify the optimal combination for model performance.
-For deep learning models, we leveraged Microsoft’s Neural Network Intelligence (NNI) 35.
-NNI is an open-source Python toolkit that automates neural architecture search and hyperparameter optimization.
-It efficiently dispatches and manages trial jobs generated by various tuning algorithms, facilitating the search for optimal neural architectures and hyperparameter configurations.
+To develop a robust predictive model for CKD progression to ESRD, we adopted methodologies from prior studies 7; 25 by constructing a longitudinal data representation with three-month intervals following initial CKD stage 3 diagnosis.
+These intervals were sequentially indexed as timestamps (0 for months 0-3, 1 for months 3-6, etc.).
+This temporal segmentation approach serves several purposes, capturing CKD’s temporal evolution, reflecting its gradual progression, aligning with clinical guidelines recommending periodic nephrologist visits, and accommodating potential time lags between clinical events and their corresponding claims data.
+These lags, particularly notable for critical events such as transplantation or dialysis initiation, exhibit substantial variability both across and within patient records, complicating consistent correction strategies.
+Employing three-month intervals mitigates the impact of these temporal discrepancies, improving model accuracy.
+Numerical features were aggregated, categorical features encoded binarily to reflect conditions or events, and baseline historical records were analyzed to adjust for pre-existing conditions, thereby ensuring accurate representation of CKD progression.
+This timestamp-based structure allows models to capture disease progression dynamics and identify critical temporal patterns signaling heightened ESRD risk.
+We evaluated multiple deep learning architectures, including Convolutional Neural Networks (CNN)26 for feature extraction; Recurrent Neural Networks (RNN)27, Long Short-Term Memory networks (LSTM)28, and Gated Recurrent Units (GRU)29 for modeling temporal dependencies; and Temporal Convolutional Networks (TCN)30 for sequence modeling.
+These architectures were selected to optimize prediction accuracy of ESRD progression and improve identification of associated risk factors.
+Detailed implementation information is available in Appendix D.
+To maximize clinical relevance, our analytical framework employs a two-phase approach (as illustrated in Fig.
+2(a) Stage 2).
+First, we systematically compare all modeling techniques mentioned above across multiple observation windows to identify optimal predictive performance.
+Second, recognizing that clinicians cannot predetermine a patient’s progression timeline to ESRD in real-world settings, we re-train the best-performing model using the complete cohort with the optimal observation window.
+This clinically-oriented model then undergoes comprehensive explainable AI analysis to identify key predictive features and temporal patterns.
+This approach bridges the gap between statistical performance and practical clinical implementation, providing actionable insights for nephrologists.
+4.3 Explainable AI Methods
+After identifying the best-performing model and optimal observation window, we employed two complementary explainable AI techniques to understand the features driving these predictions.
+At the cohort level, we utilized feature importance analysis to identify key predictive variables across the population.
+At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.
 4.4 Additional Analyses
-The comparison of these modeling results helps us identify the optimal observation length and the most reliable and accurate model for predicting CKD progression to ESRD.
-However, recognizing the limitations of focusing solely on fixed timelines, we extended our analysis to address clinical realities.
-In practice, the progression timeline from CKD to ESRD is often uncertain when a patient presents for evaluation.
-To accommodate this variability, we leveraged the selected optimal model and observation window to integrate all time-length cohorts, ranging from 6 to 30 months.
-This integration enables predictions of ESRD likelihood and optimal observation windows without requiring precise progression timelines, enhancing the model’s relevance and applicability in real-world clinical settings.
-Building on this framework, we conducted additional analyses to further refine the model, including a detailed investigation of misclassifications and their potential causes, as well as an evaluation of the impact of updated eGFR equations on racial bias in CKD predictions.
-These efforts aim to ensure fairness and accuracy in the model’s clinical applications.
+Beyond the core dual-phase predictive framework outlined above, we conducted several additional analyses to enhance the robustness and fairness of our ESRD prediction models.
+Specifically, we investigated cases of model misclassification to identify underlying patterns and contributing factors, providing deeper insights into prediction limitations.
+Furthermore, we assessed the influence of recently updated eGFR equation on racial disparities within CKD progression predictions, aiming to ensure clinical fairness and improve the accuracy of our predictive models.
 4.4.1 Analysis of Model Misclassifications
-To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set.
-This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes.
-We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error).
-By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
-This detailed error analysis is crucial for improving the model’s accuracy and ensuring its reliability in clinical settings.
-By understanding the reasons behind incorrect predictions, we can enhance the model’s ability to support clinicians in making more informed and accurate decisions regarding patient care.
+We examined predictive probability distributions across our sample set to analyze misclassifications from our optimal model.
+We evaluated both false positives (type I errors) and false negatives (type II errors) in ESRD progression predictions to identify patterns contributing to incorrect predictions.
+This error analysis provides insights into model limitations and highlights potential areas for refinement, ultimately improving clinical reliability and decision support.
 4.4.2 Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
-The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias.
-In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of its biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare.
-In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 36, which excluded the race coefficient to improve accuracy and equity 37.
-This adjustment aimed to enhance prediction precision, especially for minority populations 38; 39.
-Our analysis, using data from 2009 to 2018, applied both eGFR equations to the data.
-To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
-It is important to note that adopting the updated equation may alter the cohort used for analysis, as the new eGFR values directly influence CKD stage classification.
-This could impact the cohort identification process, and lead to the reclassification of patients, particularly those near stage boundaries.
-Such changes may affect both the composition of the cohort and the model’s predictions.
+The 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD-EPI equation that removed race coefficients, addressing concerns about racial bias in eGFR calculations 31; 32, such as race-based components lacking biological basis and risking the perpetuation of healthcare disparities 33; 34.
+For our study using data from 2009-2018, we primarily used the 2009 CKD-EPI equation since this was the standard clinical calculation during the study period and informed actual physician decisions and CKD staging.
+We additionally applied the 2021 race-free equation in our further analysis section to compare outcomes and assess improvements in prediction accuracy and equity, particularly for minority populations.
+Note that adopting the updated equation may affect CKD stage classification, potentially altering cohort composition and model predictions.
 5 Results
 5.1 Cohort for Analysis & Data Characteristics
-(a) Variation in CKD stage distribution as a function of different observation window lengths.
+(a) Variation in CKD stage distribution of merged data as a function of different observation window lengths.
 (b) Comparison of CKD stage distribution across claims, clinical, and merged data with a 30-month observation window.
 Figure 3: Overview of CKD Stage Distribution in Merged Data Across Observation Windows.
-In this subsection, we present the trends observed during our data integration process and provide a detailed overview of the data characteristics for the final cohort utilized in our model performance analysis.
-Fig.3(a) illustrates the changes in patient distribution across CKD stages over varying observation window lengths.
-As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis.
-This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period.
-Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
-To ensure model consistency and fair comparability, we trained and tested on the same cohort.
-Recognizing the challenge of shrinking cohort size with extended observation windows, we focused on patients with at least 30 months of data following CKD stage 3 diagnosis and split the test data from this group.
-For the training data, we used the same cohort but with varying data lengths corresponding to the selected observation windows.
-This approach allowed us to explore the impact of different observation windows and identify the optimal one based on the performance on test data.
-After establishing the cohort with at least 30 months of post-CKD stage 3 data, we compared the datasets derived from claims data only (5,518 patients), clinical data only (6,842 patients), and the merged dataset used for analysis.
-As shown in Fig.
-3(b), merging the datasets resulted in a substantial reduction in the number of patients—to 1,422 patients with a 24-month observation window (which proved optimal in subsequent analyses) and 1,355 patients with a 30-month observation window.
-This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis.
-Table 1: Data characteristics of the patient cohort for analysis
+In this subsection, we present data integration trends and cohort characteristics.
+Fig.3(a) shows patient distribution across CKD stages with varying observation windows.
+Patient numbers decline as windows extend from 6 to 30 months due to insufficient data or ESRD progression within the window.
+The ESRD columns represent patients who progressed to ESRD after their respective observation windows.
+While some patient losses may be attributed to mortality, we lack confirmatory data.
+To ensure model consistency and fair comparability, we trained and tested our models using the same patient cohort.
+Recognizing the issue of cohort sizes decreasing as the observation window lengthens, we selected patients who had at least 30 months of follow-up data after their CKD stage 3 diagnosis.
+We specifically chose 30 months because it represents the smallest cohort across the range of observation windows; thus, patients in this 30-month cohort are also guaranteed to be present in all shorter observation windows.
+For the training datasets, we utilized this same group of patients but varied the length of the data included according to different observation windows.
+This approach allowed us to systematically assess how different observation windows affected model performance and to identify the optimal window based on test dataset results.
+After requiring $\geq$ 30 months of post–stage 3 data, we compared claim-only (n = 5,518) and clinical-only (n = 6,842) cohorts to the merged sample.
+Combining data reduced the cohort to 1,422 patients for the 24-month window (optimal) and 1,355 for the 30-month window (Fig.
+3(b)), reflecting exclusion of those lacking one data source and yielding a more comprehensive dataset.
+Table 1: Data characteristics of the patient cohort for analysis (Merged data, 24-month observation window, n=1,422).
+Continuous variables are shown as mean ± standard deviation and compared by independent t-test (log-transformed if non-normal); categorical variables are shown as count (percentage) and compared by chi-squared test.
+“Missing data (%)” indicates the proportion of missing observations for each variable.
+P-values are two-sided, with $p<0.05$ denoting statistical significance.
 Characteristics Missing data (%) Progressed to ESRD (n = 86) Non-progressed to ESRD P-value
@@
 Number of emergency department visits 16 2.63 ± 2.18 2.01 ± 1.99 0.005
-Table 1 summarizes the demographic, comorbidity, claims-driven, and clinical-driven characteristics of the patient cohort, comprising 1,422 patients; 86 progressed to ESRD (6%), 1,336 did not.
-For patients who progressed to ESRD, the mean time from CKD stage 3 to ESRD was 4.82 ± 1.82 years.
-The distribution plot of progression times is provided in Supplementary Figure 1.
-Patients who progressed to ESRD were significantly younger ( 69.13 vs.
-72.04 years, $p<0.001$ ), with notable racial disparities.
-African American patients were more likely to progress to ESRD (14.0% vs.
-4.5%), while White patients had lower progression rates ( $81.4\%$ vs.
+Table 1 summarizes characteristics of our 1,422-patient cohort; 86 (6%) progressed to ESRD while 1,336 did not.
+Mean progression time from CKD stage 3 to ESRD was $4.82\pm 1.82$ years.
+ESRD patients were younger ($69.13$ vs.
+$72.04$ years, $p<0.001$) with significant racial disparities: African Americans showed higher progression rates (14.0% vs.
+4.5%) while White patients had lower rates (81.4% vs.
 93.0%, $p<0.001$).
-Comorbidities such as hypertension were prevalent in both groups (99%), but diabetes was more frequent in the ESRD group ( $73.3\%$ vs.
-$0.00959.0\%,p=$ 0.009 ).
-Secondary hyperparathyroidism and conduction/dysrhythmias were also significantly higher in ESRD patients ( $p<0.001$ and $p=0.005$ , respectively).
-In claims-driven features, ESRD patients showed slightly higher counts of outpatient and professional claims ( $p=0.039$ ) but no significant differences in costs.
-Clinical-driven features highlighted lower eGFR (17.21 vs.
-22.78, $p<0.001$ ), lower hemoglobin levels (12.15 vs.
-$0.00114.25,p<$ 0.001 ), and higher occurrences of advanced CKD stages in the ESRD group.
-These differences underscore the clinical and demographic factors associated with ESRD progression.
+Hypertension was prevalent in both groups (99%), but diabetes was more common in ESRD patients (73.3% vs.
+59.0%, $p=0.009$).
+Secondary hyperparathyroidism and conduction disorders were also significantly higher in the ESRD group.
+Claims data showed slightly higher outpatient/professional claims counts for ESRD patients ($p=0.039$) but no significant cost differences.
+Clinically, ESRD patients had lower eGFR ($17.21$ vs.
+$22.78$, $p<0.001$), lower hemoglobin ($12.15$ vs.
+$14.25$, $p<0.001$), and more advanced CKD stages.
 5.2 Performance Comparison of Models Across Datasets
@@
 Most notably, the merged dataset combining both claims and clinical data yielded the best overall performance, with LSTM achieving better results across all metrics (AUROC: 0.93, F1: 0.65, AUPRC: 0.61).
-Comprehensive results across all models and observation windows are provided in Supplementary Table 1.
-Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data.
+Comprehensive results across observation windows are provided in Appendix E.
+Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data (24-month observation window, n=1,422).
+Bold indicates the best performance within each method category (Machine Learning or Deep Learning).
 (a) Model performance for prediction of ESRD by using claims data.
-Model Model performance metric
+Model Type Model Model performance metric
 F1 Score AUROC AUPRC
-Logistic Regression 0.33 0.72 0.44
+Machine Learning Logistic Regression 0.33 0.72 0.44
 Random Forest 0.36 0.74 0.48
 XGBoost 0.39 0.75 0.47
-CNN 0.45 0.82 0.50
+Deep Learning CNN 0.45 0.82 0.50
 RNN 0.50 0.90 0.52
@@
 (b) Model performance for prediction of ESRD by using clinical data.
-Model Model performance metric
+Model Type Model Model performance metric
 F1 Score AUROC AUPRC
-Logistic Regression 0.54 0.76 0.47
+Machine Learning Logistic Regression 0.54 0.76 0.47
 Random Forest 0.58 0.79 0.51
 XGBoost 0.57 0.80 0.52
-CNN 0.56 0.84 0.53
+Deep Learning CNN 0.56 0.84 0.53
 RNN 0.61 0.85 0.53
@@
 (c) Model performance for prediction of ESRD by using merged data.
-Model Model performance metric
+Model Type Model Model performance metric
 F1 Score AUROC AUPRC
-Logistic Regression 0.55 0.75 0.45
+Machine Learning Logistic Regression 0.55 0.75 0.45
 Random Forest 0.60 0.84 0.49
 XGBoost 0.61 0.85 0.51
-CNN 0.56 0.80 0.46
+Deep Learning CNN 0.56 0.80 0.46
 RNN 0.62 0.87 0.53
@@
 5.3 Key Feature Analysis Using Explainable AI Techniques
-Having identified the 24-month observation window as optimal and LSTM as the best-performing model for ESRD prediction, we employed two complementary explainable AI techniques to understand the features driving these predictions.
-At the cohort level, we utilized XGBoost’s feature importance analysis to identify key predictive variables across the population.
-At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.
+Leveraging the optimal 24-month window and our best ESRD prediction model, we performed (1) cohort-level feature importance to pinpoint the most influential predictors across the population, and (2) individual-level SHAP analysis to generate patient-specific explanations of each prediction.
 5.3.1 Cohort-Level Key Feature Identification
@@
 4(a)).
-The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD.
+The most critical feature was the presence of CKD Stage 5 (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD.
 Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly.
@@
 However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
-(a) Top 30 features of importance for the XGBoost model.
+(a) Top 30 features importance for the XGBoost model. Features are colored from yellow-green (highest importance) to dark blue (lower importance).
 (b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
-Figure 4: Key Feature Analysis: Feature importance and SHAP analysis for ESRD risk prediction using a 24-month observation window.
+Red bars indicate features that increase predicted risk, while blue bars indicate features that decrease predicted risk.
+Figure 4: Feature importance and SHAP analysis for ESRD risk prediction using XGBoost model (24-month observation window, n=1,422). See Appendix G for complete feature names.
 5.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
-While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach.
-To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
-We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD.
-Despite both having CKD Stage 5, they showed notably different patterns of contributing features, highlighting how the same clinical outcome can arise from distinct combinations of risk factors and underscoring the importance of personalized treatment strategies.
-In Figure 4(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue.
-For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it.
-For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
-This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning.
-SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile and modifiable factors.
+To support personalized decision-making, we applied SHAP analysis to quantify the features driving each patient’s risk prediction.
+Figure 4(b) presents SHAP force plots for two correctly predicted high-risk patients, Patient X and Patient Y, both diagnosed with CKD Stage 5 but exhibiting distinct risk profiles.
+For Patient X, elevated risk is driven primarily by Stage 5 CKD (S5) and a high volume of outpatient claims (n_claims_DR_min), while higher eGFR levels and the absence of certain clinical markers mitigate that risk.
+In contrast, Patient Y’s lower risk contribution—despite the same Stage 5 diagnosis—stems from younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values.
 5.4 Analysis of Model Misclassifications: Type I and Type II Errors
+All subsequent analyses utilize our optimal 24-month observation window and the LSTM model, which demonstrated superior performance among all tested architectures.
 To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 5.
 Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered.
-This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
-To understand the underlying causes of model misclassification, we conducted a detailed analysis of Type I and Type II errors.
-We focused particularly on Type II errors, where patients were incorrectly predicted not to develop ESRD but ultimately did progress to the condition.
-By comparing the 16 false negatives (patients shown in the lower part of the right plot) with the 70 true positives (patients correctly predicted to progress to ESRD, shown in the upper part of the rightplot), we aimed to identify patterns and characteristics that might explain why the model failed to identify these high-risk patients.
-These errors can delay essential treatment, potentially accelerating disease progression.
-Conversely, Type I errors represent patients who were incorrectly predicted to develop ESRD but did not actually progress.
-Our analysis contrasted 17 false positives (patients shown in the upper part of the left plot) with 1,319 true negatives (patients correctly predicted not to progress, shown in the lower part of the left plot).
-These Type I errors are particularly concerning in clinical practice as they could lead to unnecessary medical interventions and inefficient allocation of healthcare resources, potentially diverting attention from patients at genuine risk of ESRD progression.
-Understanding the characteristics distinguishing these misclassified cases from correctly predicted ones may provide valuable insights for model refinement and improved clinical decision support.
+This clustering suggests that consistent factors may be influencing these errors.
 Figure 5: Analysis of Model Misclassification: Type I and Type II Errors.
-The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset.
-The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
-Table 3: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy.
+The left panel shows predicted probabilities for all patient outcomes (24-month conversation window, n=1,422), stratified by ESRD (green) and non-ESRD (blue) cases, with patient indices (1-1422) on the x-axis.
+The right panel displays only the incorrectly predicted cases (n=33), using actual patient IDs on the x-axis for identification purposes.
+The horizontal dashed line at 0.5 represents the classification threshold.
+To understand model misclassification causes, we analyzed Type I and Type II errors.
+For Type II errors, we compared 16 false negatives (patients incorrectly predicted not to develop ESRD but who did) with 70 true positives (patients correctly predicted to progress).
+For Type I errors, we contrasted 17 false positives (patients incorrectly predicted to develop ESRD) with 1,319 true negatives (patients correctly predicted not to progress).
+Table 3: Analysis of Type I and Type II Errors (24-month observation window, n=1,422): Features and Their Impact on Prediction Accuracy.
 This table presents representative subsets of features under various timestamps.
@@
 net_exp_O 1044.82 333.61 0.000324 4
-Based on Table 3(a), CKD Stage 5 emerges as a critical feature.
-For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, which is significantly lower than the 0.36 mean for those correctly predicted to progress.
-This suggests that these patients did not have a Stage 5 record until the subsequent timestamp.
-The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely.
-However, after consulting with nephrologists and examining the literature on CKD progression, we note that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
-Based on Table 3(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD.
-This is unusual.
-One possible explanation could be the issue of data censoring.
-Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later.
-Additional data collection beyond 2018 would be needed to update and refine this analysis.
+Based on Tables 3(a) and 3(b), CKD Stage 5 emerges as a critical feature in prediction errors.
+For false negatives (incorrectly predicted as not progressing to ESRD), the mean value for CKD Stage 5 at timestamp 6 is significantly lower (0.06 versus 0.36) compared to correctly predicted cases, suggesting these patients didn’t have a Stage 5 record until later.
+Conversely, false positives show CKD Stage 5 presence at both timestamps 6 and 7, indicating these patients had Stage 5 records but didn’t actually progress to ESRD.
 5.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
-(a) Incorrect predictive probabilities using the 2009 eGFR equation.
-(b) Incorrect predictive probabilities using the 2021 eGFR equation.
-(c) Demographic distribution patterns under the 2021 eGFR equation. Encoding mappings are: Race: 0=White, 1=Black, 2=Others; Gender: 0=Male, 1=Female; ESRD Status: 0=non-ESRD, 1=ESRD.
-(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
+(a) Incorrect predictive probabilities obtained using the 2009 eGFR equation.
+(b) Incorrect predictive probabilities obtained using the 2021 eGFR equation.
 Figure 6: Impact of Updated eGFR Equations (2009 vs.
-2021) on Prediction Errors and Racial Bias.
-Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
-When comparing predictions between the 2009 and 2021 eGFR equations (Fig  6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228.
+2021) on Prediction Errors.
+Subplots compare incorrect predictions using the two eGFR equations, highlighting the differences in misclassification patterns between the traditional and updated formulas.
+When comparing predictions between the 2009 and 2021 eGFR equations (Fig.
+6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228.
 Among these misclassifications, only 7 cases were common to both equations.
-Analysis of racial distribution in prediction errors (Fig  6(c),  6(d)) revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12.
-The demographic breakdown for the 2021 equation cohort, including race, gender, and ESRD outcomes, is shown in Fig  6(c).
+Analysis of racial distribution in prediction errors revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12.
 However, these numbers are clearly too small for robust insights.
@@
 This study demonstrates the value of integrating multiple data sources and deep learning methodologies for improving ESRD progression predictions in CKD patients.
-By combining claims and EHR data with LSTM and GRU models, which excel at capturing temporal dependencies, we achieved enhanced predictive accuracy using a 24-month observation window.
-This window length provides an optimal balance between early detection and prediction reliability, supporting timely clinical interventions.
-Our research addresses a critical gap in chronic disease management by integrating clinical and claims data, offering a more comprehensive view of patient health trajectories than single-source approaches.
-The combination of medical indicators and healthcare utilization patterns provides deeper insights into ESRD progression risk factors, making the model more robust across diverse patient populations.
-The determination of this optimal observation window represents a significant advancement in balancing timely interventions with prediction accuracy, enabling healthcare providers to efficiently monitor CKD progression while minimizing data collection burden.
-Our findings suggest that while the updated 2021 eGFR equation shows modest improvement in overall prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228) in this study population, the more useful insight may come from the pattern of misclassifications.
-The fact that only 7 cases were misclassified by both equations, despite similar overall error rates, indicates that these equations capture different aspects of kidney function and disease progression.
-This has important implications for clinical practice: while the 2021 equation may better serve minoritized populations by removing race-based adjustments, the persistence of prediction errors clustered in certain patient profiles suggests that eGFR alone may not be sufficient for accurate ESRD prediction.
-These findings underscore the need for a multi-faceted approach to risk assessment, combining both traditional clinical markers and novel predictive features.
-By leveraging integrated data sources and advanced modeling techniques, healthcare providers can develop more comprehensive risk assessment strategies, potentially leading to more targeted interventions and improved patient outcomes across diverse populations.
+By combining claims and EHR data with LSTM and GRU models, we achieved enhanced predictive accuracy using a 24-month observation window—an optimal balance between early detection and prediction reliability.
+Our approach addresses a critical gap in chronic disease management by offering a more comprehensive view of patient health trajectories than single-source approaches.
+Error analysis revealed important clinical implications.
+Patients incorrectly predicted not to develop ESRD (false negatives) had significantly lower prevalence of CKD Stage 5 records at timestamp 6 compared to true positives (0.06 vs 0.36), suggesting sudden kidney function decline—a phenomenon clinicians recognize as ”crashing” into dialysis.
+Conversely, false positives showing CKD Stage 5 in consecutive timestamps without progression may reflect data censoring, as patients might have progressed after our dataset endpoint.
+While the updated 2021 eGFR equation showed modest improvement in prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228), the pattern of misclassifications offers greater insight.
+With only 7 cases misclassified by both equations despite similar overall error rates, these formulas clearly capture different aspects of kidney function.
+This suggests that while the 2021 equation may better serve minoritized populations by removing race-based adjustments, eGFR alone remains insufficient for accurate ESRD prediction.
+These findings underscore the need for a multi-faceted approach to risk assessment, combining traditional clinical markers, novel predictive features, and awareness of potential rapid disease progression patterns to develop more targeted interventions across diverse populations.
 6.1 Limitations
-Our study’s reliance on data from a single institution limits the model’s generalizability to all care delivery settings.
-Incorporating data from more sources could enhance robustness and applicability across different clinical settings.
-The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
-The use of EHR data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness.
-Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
-Despite employing data oversampling methods, the disparity between our model’s high AUROC scores and relatively lower F1 and AUPRC scores indicates persistent challenges in addressing data imbalance.
-This gap suggests that traditional oversampling techniques may not fully compensate for the inherent imbalance in ESRD progression data.
-To enhance performance, exploring advanced data balancing methods is essential.
-Techniques such as ensemble methods, cost-sensitive learning, and hybrid approaches that combine oversampling and undersampling have shown promise in effectively managing imbalanced datasets.
-Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions.
-Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
+This study’s reliance on data from a single institution may limit the model’s generalizability to other care settings.
+Using EHR data introduces observational bias, incomplete records, and underrepresentation of certain patient groups, which can undermine both accuracy and fairness.
+Although we applied oversampling to mitigate class imbalance, the contrast between high AUROC and lower F1/AUPRC indicates that imbalance remains an issue; more sophisticated approaches—such as ensemble methods, cost-sensitive learning, or hybrid sampling—may be needed.
+Finally, unaddressed time lags between claims and clinical data can distort temporal relationships and reduce predictive precision.
 6.2 Future Directions
-Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018.
-Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial.
-This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
-Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats.
-This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
-Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
-Expansion to Other Chronic Diseases: To evaluate framework’s versatility, we plan to apply it to other chronic conditions, such as heart disease.
-This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
+To address data censoring issues, we will first truncate data to 2016 and analyze outcomes from 2017-2018, then expand our dataset beyond 2018 to enhance trajectory modeling.
+We’ll integrate unstructured clinical notes to capture patient information missed in structured data, providing a more comprehensive health view and improving prediction accuracy.
+We will implement advanced algorithms to synchronize claims and clinical data temporally, reducing time lag effects that currently impact model performance.
+To understand prediction errors, we’ll perform SHAP analysis on misclassified cases, identifying key features contributing to these misclassifications and guiding targeted model improvements.
+Finally, we’ll validate our framework’s versatility by applying it to other chronic conditions such as heart disease, assessing its broader potential across various care delivery settings.
+This systematic expansion will provide actionable insights for both our ESRD prediction model and chronic disease management more broadly.
 7 Conclusion
-As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods.
-Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression.
-Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
-A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making.
-The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions.
-Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
-Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management.
-These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
+This study demonstrates the effectiveness of integrating diverse healthcare data with advanced machine learning techniques to accurately predict ESRD in CKD patients.
+The combined-data approach substantially enhances predictive performance and provides deeper insights into disease progression.
+SHAP analysis and feature importance assessment highlighted key predictors at both individual and cohort levels.
+A critical contribution of this work is a framework for optimizing observation windows, balancing early detection with prediction accuracy.
+We identified a 24-month observation window as optimal for maximizing predictive effectiveness while minimizing unnecessary interventions.
+Additionally, evaluating the updated 2021 eGFR equation supports efforts towards health equity and fairer clinical outcomes.
+Overall, this research advances CKD management through integrated data and innovative AI methodologies, setting the stage for personalized, equitable care and future applications in chronic disease prediction and management.
 8 Acknowledgments
@@
 Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice and for the invaluable help provided by their data team in understanding the claims data.
-AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance.
-GPT-4o and Claude 3.5 (Sonnet) were used solely for polishing grammar and refining expression.
-The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity.
-Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
 Funding Statement
@@
 The authors have no competing interests to declare.
+Contributorship Statement
+Both authors meet the International Committee of Medical Journal Editors (ICMJE) authorship criteria.
+Rema supervised the research, advised on study conception and design, reviewed and critically revised the manuscript.
+Yubo contributed to study conception and design, implemented all computational methods, wrote the manuscript draft, and performed data analysis.
+Both authors approved the final manuscript version and agreed to be accountable for all aspects of the work, ensuring its accuracy and integrity.
+AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance.
+GPT-4.5 and Claude 3.7 (Sonnet) were used solely for polishing grammar and refining expression.
+The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity.
+Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
 Data Availability Statement
@@
 Supplementary
-Supplementary Table 1: Model Performance Metrics Across Observation Windows
-Table 4: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data.
+[…]
+Appendix A Detailed Data Source Information
+Administrative Claims Data
+The de-identified administrative claims dataset used in this study was provided by Highmark, a major health insurance provider serving Pennsylvania, Delaware, and West Virginia.
+The dataset includes comprehensive records of patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018.
+The use of this dataset was approved by Highmark’s Institutional Review Board.
+Clinical EHR Data
+The clinical dataset was sourced from the Electronic Health Record (EHR) data of The Nephrology Associates (TMA, P.C.), a leading community nephrology practice serving the greater Pittsburgh region.
+This dataset contains detailed laboratory results, patient demographics, diagnostic information, and medication records spanning from 1998 to 2018.
+For consistency of integration with the claims data, we truncated the clinical data to match the 10-year span of the claims data (2009–2018).
+The usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
+Appendix B Laboratory Test Value Ranges for Distribution Detection and Adjustment
+Table 4: The table below presents the reference ranges established for laboratory tests during the data preprocessing phase.
+These ranges were used for distribution detection and adjustment to identify and correct unit inconsistencies.
+Laboratory Test Minimum Value Maximum Value Unit Conversion Notes
+Serum Creatinine 0 15 $1\mathrm{mg}/\mathrm{dL}=88.4\mu\mathrm{~{}mol}/\mathrm{L}$
+Intact PTH 0 500 $50\mathrm{pg}/\mathrm{mL}=5.263\mathrm{pmol}/\mathrm{L}$
+Hemoglobin 2 20 $6451\mathrm{~{}g}/\mathrm{dL}=10\mathrm{~{}g}/\mathrm{L};1\mathrm{~{}g}/\mathrm{dL}=100/$ 645 $\mathrm{mmol}/\mathrm{L}$
+Albumin to Creatinine Ratio, Urine 0 600 $1\mathrm{mg}/\mathrm{g}=1000\mathrm{mg}/\mathrm{mg}$
+Bicarbonate (Mostly Arterial) 0 30 $\mathrm{mEq}/\mathrm{L}=\mathrm{mmol}/\mathrm{L}$
+Phosphorus 0 10 $4\mathrm{mg}/\mathrm{dL}=0.00129\mathrm{mmol}/\mathrm{L}$
+Serum Calcium 0 20 $10\mathrm{mg}/\mathrm{dL}=0.00250\mathrm{mmol}/\mathrm{L}$
+Urine Albumin 0 600 -
+eGFR 0 150 -
+Appendix C Comorbidity Feature ICD Code Prefixes
+Table 5: Comorbidity features with ICD-9 and ICD-10 code prefixes used for feature extraction in this study
+Condition ICD-9 Prefixes ICD-10 Prefixes
+Diabetes 249*, 250* E08*–E13*
+Hypertension 401*–405* I10*–I15*
+Cardiovascular Disease 429* I25*
+Anemia 285.9 D64.9, D50*–D59*
+Metabolic Acidosis 276.2 E87.2
+Proteinuria 791.0 R80.9
+Secondary Hyperparathyroidism 252.02, 588.81 E21.1, N25.81
+Phosphatemia 275.3 E83.39
+Atherosclerosis 440* I70*
+Congestive Heart Failure (CHF) 428* I50*
+Conduction & Dysrhythmias 426*–427* I44*–I49*, Z450*, Z958*
+Myocardial Infarction (MI) 410* I21*
+CVD Other 390*–459* I05*–I09*, I30*–I52*
+Stroke 433*–434* I63*
+Fluid/Electrolytes 276* E86*–E87*
+Mineral Disorders 275* E20*–E21*, E83*, N25*
+Nutritional Deficiencies 260*–269* E40*–E46*, E50*–E64*, D50*
+CKD Stage 4 585.4* N18.4*
+CKD Stage 5 585.5* N18.5*
+CKD End-Stage Renal Disease (ESRD) 585.6* N18.6*
+Note: prefix * matches any sequence of trailing digits. For example, I44*–I49* captures all codes from I440 through I499 such as I4402 and I4439; Z450* captures Z4502, Z45010, Z45018, etc.
+Appendix D Neural Network Implementation Details
+Hardware Specifications
+Resource Specification
+GPU 2 $\times$ NVIDIA RTX 4090 (24GB VRAM)
+RAM 512GB DDR4
+Storage 2TB NVMe SSD
+Table 6: Hardware specifications used for model training and evaluation.
+Software Environment and Library Versions
+Component Version Purpose
+Python 3.10.8 Programming language
+CUDA 11.6 GPU acceleration library
+PyTorch 1.13.0 Deep learning framework
+H2O 3.38.0.1 Machine learning model tuning
+NNI 2.10 Neural architecture search
+scikit-learn 1.1.3 Traditional ML algorithms
+pandas 1.5.1 Data manipulation and analysis
+NumPy 1.23.4 Numerical computing
+Table 7: Software environment and library versions used for model implementation.
+Model Architecture, Hyperparameters, and Training Details
+Model Architecture Details Parameters Hyperparameters Training Time
+CNN 3 conv layers (64, 128, 256 filters), 2 FC layers (512, 256) 4.3M Batch size: 32, Optimizer: Adam, Learning rate: 1e-4, Dropout: 0.3 8.5 hours
+RNN 2 recurrent layers (128 units), 2 FC layers (256, 128) 2.1M Batch size: 64, Optimizer: RMSprop Learning rate: 5e-4, Dropout: 0.2 9.2 hours
+LSTM 2 LSTM layers (256 units), 1 FC layer (128) 5.7M Batch size: 32, Optimizer: Adam Learning rate: 1e-4, Dropout: 0.3 15.3 hours
+GRU 2 GRU layers (256 units), 1 FC layer (128) 4.9M Batch size: 32, Optimizer: Adam Learning rate: 1e-4, Dropout: 0.3 13.7 hours
+TCN 4 temporal blocks, kernel size 3, 128 filters 3.8M Batch size: 32, Optimizer: Adam Learning rate: 2e-4, Dropout: 0.2 12.1 hours
+Table 8: Model architecture, hyperparameters, and training details for each neural network model.
+Appendix E Model Performance Metrics Across Observation Windows
+Table 9: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data (24-month observation window, n=1,422).
 Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
@@
 TCN 0.71 0.49 0.74 0.57 0.88 0.60 0.89 0.61 0.91 0.59
-Supplementary Figure 1
-Figure 7: Distribution of time to ESRD for the ESRD cohort.
-The plot illustrates the variability in time to ESRD across the cohort, with a mean time of 4.82 years and a standard deviation of 1.82 years.
-Supplementary Table 2
-Table 5: Mapping between Feature Abbreviations in The Feature Importance Plot and Full Names
+Appendix F Distribution of time to ESRD for the ESRD cohort
+Figure 7: The plot illustrates the variability in time to ESRD across the cohort, with a mean time of 4.82 years and a standard deviation of 1.82 years.
+Appendix G Feature Abbreviations
+Table 10: Mapping between feature abbreviations in The feature importance plot and corresponding full names.
 Short Name Full Name
-S5 Serum Albumin
+S5 CKD Stage 5
 MA Metabolic Acidosis